### v2

# Arbitrary Entropy Policy Optimization: Entropy Is Controllable in Reinforcement Finetuning

Abstract
Reinforcement finetuning (RFT) is essential for enhancing the reasoning capabilities of large language models (LLM), yet the widely adopted Group Relative Policy Optimization (GRPO) suffers from entropy collapse, where entropy monotonically decreases, exploration vanishes, and policies converge prematurely. Existing entropy-regularized methods only partially alleviate this issue while introducing bias and instability, leaving entropy control unresolved and the connection between entropy, exploration, and performance unclear. We propose Arbitrary Entropy Policy Optimization (AEPO), which eliminates entropy collapse by replacing entropy bonuses with REINFORCE policy gradient on temperature-adjusted distributions and stabilizing entropy through temperature regulation. AEPO integrates three key designs: policy gradient as regularization, distribution as regularization, and REINFORCE as regularization, enabling precise entropy control without distorting optimization. Experiments demonstrate three major contributions: AEPO (1) stabilizes entropy at arbitrary target levels, effectively removing collapse in GRPO; (2) reveals a non-monotonic relation where performance first improves then declines with increasing entropy, clarifying the link between entropy, exploration, and reasoning; and (3) generalizes beyond entropy, providing a broader RFT paradigm where superior target distributions can serve as REINFORCE regularizers.
Machine Learning, ICML
[…]
inkscapearea=page
Figure 1: Entropy across five runs of AEPO. By adjusting only the parameter $\mathcal{H}$, entropy can be stably maintained at different levels.
1 Introduction
Reinforcement finetuning (RFT) has become a cornerstone for enhancing the reasoning capabilities of large language models (LLM) (GLM et al., 2024; Touvron et al., 2023; Schulman et al., 2017; Rafailov et al., 2023; Zhong et al., 2024; Wang et al., 2024). Among existing approaches, Group Relative Policy Optimization (GRPO) has gained wide adoption due to its efficiency and scalability (Shao et al., 2024; Liu et al., 2024; Guo et al., 2025). However, GRPO suffers from a well-documented drawback: entropy collapse: As training progresses, policy entropy declines monotonically, sampled outputs converge to nearly identical solutions, and the model prematurely adopts a deterministic policy with limited exploration (Yu et al., 2025; Li et al., 2025a; Zhang et al., 2025). This severely restricts the ability of RFT to discover diverse reasoning strategies. Existing remedies typically rely on entropy regularization and clip refine, which inevitably introduce bias and instability into the optimization objective, creating a trade-off between exploration and stability (Hou et al., 2025; Cui et al., 2025; Cheng et al., 2025; Shen, 2025). EFRame is the first to achieve large-scale entropy control, but it remains a qualitative analysis without conducting a quantitative study of entropy (Wang et al., 2025a).
Although the issue of entropy collapse in GRPO has been repeatedly noted, it remains unresolved: there is still no principled algorithm capable of precisely regulating entropy throughout training. Existing approaches often treat entropy only as a side indicator—for instance, using it to split steps or guide auxiliary heuristics—while the policy entropy itself continues to collapse as training progresses (Wang et al., 2025b; Zheng et al., 2025; Li et al., 2025b; Cheng et al., 2025). As a result, the relationship among entropy, exploration, and performance remains unclear: it is uncertain whether entropy is a sufficient proxy for exploration or whether exploration itself consistently improves training outcomes. If significant performance improvement can be observed by adjusting entropy to a better range, it would indicate that exploration plays a crucial role in this process. If arbitrary entropy control can be achieved during RFT, it would make it possible to realize exploration at any desired degree and, in turn, establish a principled connection among entropy, exploration, and performance.
Motivated by these challenges, we propose Arbitrary Entropy Policy Optimization (AEPO), a novel policy gradient that directly addresses entropy collapse. AEPO replaces conventional entropy bonuses with a REINFORCE policy gradient (Williams, 1992; Sutton et al., 1999) applied to samples drawn from temperature-adjusted distributions. Entropy is further stabilized through temperature regulation, ensuring that batch-level entropy remains oscillating around an arbitrary constant $\mathcal{H}$ throughout training, as shown in Fig. 1. AEPO achieves entropy control through three key design components:
  • Policy gradient as regularization: Instead of relying on a conventional entropy bonus, we employ a full policy gradient term applied to samples that naturally exhibit high- or low-entropy properties. This design ensures that entropy never dominates the optimization objective, allowing the model to monotonically explore toward higher accuracy throughout training.
  • Distribution as regularization: Entropy is regulated via temperature-adjusted distributions. Given a predefined entropy threshold $\mathcal{H}$, when the previous step’s entropy $\mathcal{H}(\pi_{\theta_{\text{old}}})<\mathcal{H}$, we regard the distribution under $T_{\text{high}}>1$ as a better candidate and mix in a small proportion of its samples. Conversely, when $\mathcal{H}(\pi_{\theta_{\text{old}}})\geq\mathcal{H}$, we instead sample from the distribution under $T_{\text{low}}<1$.
  • REINFORCE as regularization: In Reinforcement Learning with Verifiable Reward (RLVR), the REINFORCE algorithm can filter out negative samples in an unbiased manner, allowing positive samples to form a unidirectional gradient toward a better distribution. This guides the model to optimize from positive samples that align with the target distribution. If negative samples were included at this stage, the entropy control mechanism would fail.
In summary, our contributions are threefold:
  • Controllable entropy: We propose Arbitrary Entropy Policy Optimization that can stabilize entropy at arbitrary target levels, effectively eliminating entropy collapse in GRPO.
  • Entropy–performance relation: We find that merely adjusting entropy can directly influence training performance, providing explicit evidence for the correlation between entropy, exploration, and performance. Moreover, we observe a non-monotonic trend in which performance first increases and then decreases as entropy grows, highlighting the existence of an optimal entropy regime.
  • Generalizability: Beyond entropy control, AEPO provides a broader paradigm for RFT. When a target distribution $\pi^{*}$ is identified as superior to the current policy $\pi_{\theta}$, samples from $\pi^{*}$ can be used to construct a REINFORCE policy gradient as a regularization term, which allows $\pi$ to progressively approximate $\pi^{*}$ during long-horizon training.
2 Related Work
RFT has become a central paradigm for post-training large language models (LLMs), aligning them with human feedback and task-specific objectives. Early methods such as RLHF leveraged policy optimization (e.g., PPO) to encode human preferences (OpenAI, 2023; Team et al., 2024; Wei et al., 2023; Liu et al., 2023), while Direct Preference Optimization (DPO) (Rafailov et al., 2024) later improved efficiency by optimizing policies directly from preference data. Recent models like DeepSeek-R1 (Guo et al., 2025) and Kimi-1.5 (Team et al., 2025) extend RFT through hybrid reward formulations and scalable optimization. Among them, Group Relative Policy Optimization (GRPO) (Shao et al., 2024; Liu et al., 2024) has become the de facto baseline for reasoning-focused RFT, yet suffers from entropy collapse that limits exploration of diverse reasoning strategies.
Entropy has long been regarded as a proxy for exploration in reinforcement optimization. Classical methods employ entropy regularization to stabilize training and encourage diversity (Sutton et al., 1999; Williams, 1992). More recent studies extend this idea to GRPO by introducing entropy bonuses into rewards or advantages (Cheng et al., 2025; Cui et al., 2025; Shen, 2025). However, such approaches only yield coarse-grained effects: entropy still collapses as training proceeds, or the added bias destabilizes optimization. EFRame (Wang et al., 2025a) represents the first framework to achieve large-scale entropy control through conditional sampling and replay, yet its treatment remains qualitative and lacks quantitative characterization of the entropy–exploration–performance relationship.
In summary, existing methods lack a principled mechanism to precisely regulate entropy throughout training. Moreover, the role of entropy in driving exploration and its connection to downstream performance has not been quantitatively established. Our work addresses this gap by proposing Arbitrary Entropy Policy Optimization (AEPO), which enables controllable entropy regulation and provides explicit evidence of a non-monotonic relationship between entropy, exploration, and reasoning performance.
3 Preliminary
Our work focuses on fine-tuning LLM using Reinforcement Learning (RL) for tasks with verifiable solutions, such as mathematical reasoning and code generation.
Suppose the LLM is a softmax policy, that is
$\pi_{\theta}(o_{t}|q_{t})=\frac{\text{exp}(l(q_{t},o_{t}))}{\sum_{o^{\prime}_{t}}\text{exp}(l(q_{t},o^{\prime}_{t}))},$
where $q_{t}$ is the concatenation of query q followed by $o_{<t}$, and $l(q_{t},o_{t})$ is the logit of token $o_{t}$ given input $q_{t}$. Furthermore, given a temperature T, we define:
$\pi_{\theta}^{T}(o_{t}|q_{t})=\frac{\text{exp}(l(q_{t},o_{t})/T)}{\sum_{o^{\prime}_{t}}\text{exp}(l(q_{t},o^{\prime}_{t})T)}.$
3.1 Policy-gradient based RL algorithms
Given a query $q$, let $o$ denote a response sampled from policy $\pi_{\theta}$ for query $q$. Given a reward function:
$R(q,o)=\textbf{1}[o=o^{*}],$
where $o^{*}$ is the reference response for query $q$, the policy objective is:
$\mathcal{J}(\theta)=\mathbb{E}_{q\sim P(Q),o\sim\pi_{\theta}(O|q)}\sum_{t=1}^{|o|}[R(q,o)].$
To optimize the objective function, it is a common practice to use the Policy Gradient algorithm for gradient estimation:
$\begin{split}&\nabla_{\theta}\mathcal{J}_{REINFORCE}(\theta)=\mathbb{E}_{q\sim P(Q),o\sim\pi_{\theta}(O|q)}\\ &\quad\quad\quad\quad\quad\quad\sum_{t=1}^{|o|}[\nabla_{\theta}\text{log}\pi_{\theta}(o_{t}|q,o_{<t})\cdot R(q,o)],\\ &\nabla_{\theta}\mathcal{J}_{GRPO}(\theta)=\mathbb{E}_{q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta}(O|q)}\\ &\quad\quad\quad\quad\quad\quad\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}[\nabla_{\theta}\text{log}\pi_{\theta}(o_{t}|q,o_{<t})\cdot\hat{A}_{i,t}],\end{split}$
where $R(q,o)$ is the reward for a query–response pair $(q,o)$, and $\hat{A}_{i,t}$ denotes the estimated advantage. To reduce gradient variance, GRPO extends REINFORCE by introducing group-wise normalization: for each query $q$, it samples $G$ responses $\{o_{i}\}_{i=1}^{G}$ and normalizes their rewards to compute the relative advantage for stable optimization.
$\hat{A}_{i,t}=\frac{R(q,o_{i})-\text{mean}(\{R(q,o_{j})\}_{j=1}^{G})}{\text{std}(\{R(q,o_{j})\}_{j=1}^{G})}.$
3.2 Entropy-regularization variants
In traditional RL, it is common to add an entropy term to the objective to prevent the policy from becoming overly deterministic. Prior work has also explored various approaches in this direction for LLM training.
Entropy-Reg  Given a query $q$, let $o$ denote a response sampled from policy model $\pi_{\theta}$ for query $q$. For each token $o_{t}$ in response $o$, we denote the token-level entropy as:
$\mathcal{H}_{t}(\pi_{\theta}):=-\mathbb{E}_{o_{t}\sim\pi_{\theta}(\cdot|q,o_{<t})}[\text{log}\pi_{\theta}(o_{t}|q,o_{<t})],$
and then we can further denote that:
$\mathcal{H}(\pi_{\theta}):=\mathbb{E}_{q\sim P(Q),o\sim\pi_{\theta}(O|q)}\frac{1}{|o|}\sum_{t=1}^{|o|}\mathcal{H}_{t}(\pi_{\theta}).$
In maximum entropy RL, we optimize for the entropy-regularized objective as follows:
$\mathcal{J}$ $_{MaxEnt}(\theta)=\mathcal{J}(\theta)+\lambda\cdot\mathcal{H}(\pi_{\theta})=$
$\mathbb{E}_{q\sim P(Q),o\sim\pi_{\theta}(O|q)}\sum_{t=1}^{|o|}[R(q,o)-\lambda\cdot\text{log}\pi_{\theta}(o_{t}|q,o_{<t})].$
Entropy-Adv(Cheng et al., 2025) proposed an entropy-guided advantage shaping method. The key idea is to inject an entropy-based term into the advantage function during policy optimization. They define an entropy-based advantage term $\psi(\mathcal{H}_{t})$ and use it to shape advantage:
$\psi(\mathcal{H}_{t})=\text{min}(\beta\cdot\mathcal{H}_{t}^{detach},\frac{|\hat{A}_{i,t}|}{\kappa}),$
$A_{i,t}^{\text{shaped}}=\hat{A}_{i,t}+\psi(\mathcal{H}_{t}),$
where $\beta>0\ and\ \kappa>1$. The entropy term $\mathcal{H}_{t}^{detach}$ is detached from the computational graph during backpropagation, acting as a fixed offset to the original advantage. The policy gradient of the algorithm retains a format similar to that in GRPO, where only the advantage $\hat{A}_{i,t}$ is replaced by the shaped one:
$\nabla_{\theta}\mathcal{J}^{\text{shaped}}(\theta)=$ $\mathbb{E}_{q\sim P(Q),o\sim\pi_{\theta}(O|q)},$
$\sum_{t=1}^{|o|}[\nabla_{\theta}\text{log}\pi_{\theta}(o_{t}|q,o_{<t})\cdot A_{i,t}^{\text{shaped}}].$
Table 1: Comparison of entropy-based optimization objectives.
Policy Gradient
Entropy-Reg $\sum_{t}(\hat{A}_{t}\cdot\nabla_{\theta}\text{log}\pi_{\theta}(o_{t}|q,o_{<t})+\lambda\cdot\nabla_{\theta}\mathcal{H}_{t})$
Entropy-Adv $\sum_{t}A_{i,t}^{\text{shaped}}\cdot\nabla_{\theta}\text{log}\pi_{\theta}(o_{t}|q,o_{<t})$
4 Method
Arbitrary Entropy Policy Optimization (AEPO) is designed to achieve precise and stable control of policy entropy during RFT. Unlike conventional entropy-regularized methods, AEPO does not introduce explicit entropy bonuses; instead, it adjusts the training distribution and policy gradient to regulate entropy implicitly yet controllably. This section first introduces the key premises underlying AEPO’s design, which reveal the relationship between temperature, entropy, and policy updates, and then details the algorithmic formulation of AEPO.
$\mathcal{J}_{GRPO}(\theta)=\;\mathbb{E}_{q\sim P(Q),\;\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O|q)}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\Big[r_{i,t}(\theta)\,\hat{A}_{i,t},\;\text{clip}\!\big(r_{i,t}(\theta),1-\epsilon,1+\epsilon\big)\,\hat{A}_{i,t}\Big],$ (1)
$\mathcal{J}_{AEPO}(\theta)=\;\mathcal{J}_{GRPO}(\theta)+\alpha\;\mathbb{E}_{q\sim P(Q),\;\{o_{i}\}_{i=1}^{G}\sim\pi^{T}_{\theta_{\text{old}}}(O|q)}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\Big[r_{i,t}(\theta)\,R(q,o_{i}),\;\text{clip}\!\big(r_{i,t}(\theta),1-\epsilon,1+\epsilon\big)\,R(q,o_{i})\Big],$
where $r_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}\mid q)}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid q)}$ and $T=T_{\text{low}}+\big(T_{\text{high}}-T_{\text{low}}\big)\,\mathbf{1}\!\left[\,\mathcal{H}(\pi_{\theta_{\text{old}}})<\mathcal{H}\,\right]$.
4.1 Premise
The design of AEPO is built upon two empirical premises that connect temperature-based sampling to entropy dynamics. These premises establish the foundation for controllable entropy modulation without introducing bias into the optimization objective.
Premise 1.
Higher temperature distributions globally correspond to higher policy entropy, while lower temperature corresponds to lower entropy.
Previous studies, such as Du et al. (2025), show that increasing the sampling temperature generally broadens the model’s output distribution and raises its entropy. Although the correspondence between higher temperature and higher entropy is not a universally proven principle, it holds—or approximately holds—under most intuitive and commonly used model settings. In the context of RFT, temperature thus serves as a practical external variable for indirectly regulating entropy, enabling controllable entropy adjustment.
Figure 2: Entropy dynamics under temperature-controlled sampling. High-temperature positive samples increase entropy, promoting exploration, while low-temperature positive samples reduce entropy, leading to more deterministic behavior.
Lemma 4.1.
(Cui et al., 2025) If the actor policy $\pi_{\theta}$ is a tabular softmax policy updated via natural policy gradient (Kakade, 2001) with step size $\eta$. Then the change in policy entropy between two steps approximately satisfies:
$\begin{split}\mathcal{H}(\pi_{\theta}^{k+1})-\mathcal{H}(\pi_{\theta}^{k})\approx&-\eta\cdot\mathbb{E}_{s\sim d_{\mu}^{k}}\text{Cov}_{a\sim\pi^{k}_{\theta}(\cdot\mid s)}\\ &\!\bigl[\log\!\pi^{k}_{\theta}(a\mid s),\,A^{\pi^{k}}(s,a)\bigr].\end{split}$
The proof can be seen in Liu (2025) and Cui et al. (2025). $\mathcal{H}$ indicates the policy entropy of policy model, and Cov denotes covariance, $\pi^{k}_{\theta}$ is the policy at step $k$, and $A^{\pi^{k}}(s,a)$ is the advantage function of action $a$ under state $s$. This result indicates that when positive actions have a low probability. If the probability of sampling low-probability actions can be increased, the model will be optimized toward higher entropy.
Premise 2.
Positive samples from temperature-adjusted distributions induce predictable entropy change during training.
When positive samples are drawn from a higher-temperature version of the current policy distribution, training with these samples increases the model’s entropy, promoting exploration. Conversely, sampling from a lower-temperature distribution produces entropy reduction, guiding the model toward more deterministic behavior. Figure 2 illustrates this phenomenon: the entropy increases under high-temperature sampling and decreases under low-temperature sampling. Meanwhile, entropy increase is typically slower than entropy decrease, which aligns with intuition. This dynamic forms the empirical basis for AEPO’s temperature-controlled entropy feedback loop, enabling bidirectional regulation of entropy around a target value.
4.2 AEPO
Building upon the premises introduced above, AEPO is designed to achieve stable and controllable entropy regulation throughout RFT. Its core idea is to replace explicit entropy bonuses with a REINFORCE regularization term that adjusts the sampling distribution according to the current entropy state. In this way, AEPO modulates exploration implicitly—through data distribution—rather than directly through the loss function.
Formally, AEPO augments the standard GRPO objective with an additional policy gradient term applied to samples drawn from temperature-adjusted distributions. The overall policy gradient can be expressed as:
$\nabla_{\theta}\mathcal{J}_{AEPO}(\theta)$
$=$ $\underbrace{\mathbb{E}_{q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta}(O|q)}\sum_{t=1}^{|o|}[\nabla_{\theta}\text{log}\pi_{\theta}(o_{t}|q,o_{<t})\cdot\hat{A}_{i,t}]}_{\text{GRPO form policy gradient}}$
$+$ $\underbrace{\alpha\cdot\mathbb{E}_{q\sim P(Q),o\sim{\color[rgb]{0.828125,0.18359375,0.18359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.828125,0.18359375,0.18359375}\pi_{\theta}^{T}(O|q)}}\sum_{t=1}^{|o|}[\nabla_{\theta}\text{log}\pi_{\theta}(o_{t}|q,o_{<t})\cdot{\color[rgb]{0.828125,0.18359375,0.18359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.828125,0.18359375,0.18359375}R(q,o)}]}_{\text{REINFORCE form policy gradient}}.$
The implementation of AEPO’s loss function is shown in Eq. (1), which consists of three key design components:
Policy gradient as regularization. AEPO replaces conventional entropy regularization with a full policy gradient term that simultaneously enables entropy control and prevents entropy from dominating the optimization process. As illustrated in Fig. 3, entropy regularization alone tends to drive the optimization toward two extremes—either entropy collapse or entropy explosion. In the former case, the regularization term is too weak to reverse the monotonic entropy decay; in the latter, entropy becomes an irreversible dominant factor in optimization. By contrast, AEPO constrains entropy fluctuation within a narrow and stable range through the policy gradient mechanism, making it remarkably robust to hyperparameters.
Figure 3: Comparison between entropy regularization and AEPO. Entropy regularization often drives optimization toward two extremes—collapse or explosion—while AEPO maintains entropy within a stable and optimal exploration range.
Distribution as regularization. AEPO regulates the optimization direction by adjusting the expected sampling distribution $\pi^{*}$ derived from the current policy. When the observed entropy $\mathcal{H}(\pi_{\theta_{\text{old}}})$ is below the target threshold $\mathcal{H}$, AEPO samples from the higher-temperature distribution $\pi^{T_{\text{high}}}_{\text{old}}$ to encourage exploration. Conversely, when $\mathcal{H}(\pi_{\theta_{\text{old}}})$ exceeds $\mathcal{H}$, AEPO samples from the lower-temperature distribution $\pi^{T_{\text{low}}}_{\text{old}}$ to promote stability. This bidirectional regulation mechanism achieves fine-grained entropy control, allowing the policy to maintain equilibrium between exploration and convergence.
REINFORCE as regularization. In RLVR, the reward space is binary. This property allows REINFORCE to filter out negative samples in an unbiased manner, ensuring that the gradient is formed from positive samples that align with the desired distribution. Consequently, AEPO’s regularization term produces a one-sided optimization signal that guides the policy toward higher-quality behavior distributions.
Table 2: Main results on mathematical reasoning benchmarks. We compare Qwen2.5-math-7B (base model), GRPO, entropy-based baselines, and AEPO under different entropy thresholds $\mathcal{H}$. AEPO consistently outperforms all baselines, achieving stable entropy control and delivering substantial improvements across diverse benchmarks, with the best overall performance at $\mathcal{H}=0.75$.
Benchmarks AIME24 AMC Challenge math GSM8K MATH Minerva math Olympiad Average
Qwen2.5-math-7B 13.3 40.0 41.7 65.4 65.5 11.0 26.7 37.66
GRPO 36.7 75.0 48.2 88.9 80.5 34.6 41.8 57.96
Entropy-Reg 36.7 75.0 47.6 87.0 80.4 35.7 40.4 57.39
Entropy-Adv 36.7 75.0 47.8 87.8 80.4 37.5 42.1 58.18
AEPO $\mathcal{H}=0.25$ 40.0 77.5 48.5 89.4 80.6 33.5 42.2 58.81
AEPO $\mathcal{H}=0.50$ 43.3 82.5 48.8 89.5 81.6 38.2 43.0 60.99
AEPO $\mathcal{H}=0.75$ 50.0 (+36.7) 80.0 (+40.0) 48.2 (+6.50) 89.4 (+24.0) 82.0 (+16.5) 37.5 (+26.5) 42.4 (+15.7) 61.36 (+23.70)
AEPO $\mathcal{H}=1.00$ 33.3 75.0 48.4 88.7 80.8 37.9 42.1 58.03
5 Experiments
To validate the effectiveness of our methods, we present the experimental setup and results in the following sections. Our work is based on the EasyR1 and VeRL frameworks (Yaowei et al., 2025; Sheng et al., 2025), and we compare with the RL baselines GRPO (Shao et al., 2024) and its entropy-regularization variants (Hou et al., 2025; Cheng et al., 2025).
5.1 Experimental setup
Model and Dataset: We conduct experiments to evaluate the effectiveness of AEPO in RFT for mathematical reasoning tasks. The base model is Qwen2.5-Math-7B (Yang et al., 2024), an LLM specialized for mathematical problem solving. For training, we use the DAPO-17K dataset (Yu et al., 2025), which contains diverse problem instances curated for RFT.
Benchmark: Evaluation is performed on a broad suite of mathematical reasoning benchmarks, including AIME24 (HuggingFaceH4, 2025), AMC (Lightman et al., 2023), College Math (Zhong et al., 2023), GSM8K (Cobbe et al., 2021), MATH (Lightman et al., 2023), Minerva Math (Lewkowycz et al., 2022), and Olympiad (Lightman et al., 2023). These benchmarks collectively span a wide range of difficulty levels, from grade school arithmetic to advanced competition-level mathematics, and together they cover nearly all mainstream benchmarks for mathematical reasoning, enabling a comprehensive assessment of AEPO’s impact on reasoning performance across diverse tasks.
Implementation: We initialize all policy models and conduct experiments on 8 A800 GPUs (40GB). Unless otherwise specified, we follow the default EasyR1 settings: maximum response length of 2048, global batch size 128, rollout batch size 512, rollout group size GG=5, temperature 1.0, learning rate $10^{-6}$, $\epsilon=0.2$, and a binary reward. For AEPO, $T_{\text{high}}=1.2$, $T_{\text{low}}=0.8$, we replace 60 temperature-adjusted positive samples in each batch. For entropy-regularized baselines, we set $\lambda=0.03$ for Entropy-Reg, and $\beta=0.4$, $\kappa=2$ for Entropy-Adv (Cheng et al., 2025). To exclude potential confounding factors, we did not apply KL divergence in any of our experiments.
5.2 Main results
Table 2 summarizes the performance of AEPO compared with GRPO and entropy-based baselines across seven mathematical reasoning benchmarks. We observe that AEPO consistently outperforms all competing methods, achieving the best results on every benchmark. In particular, AEPO yields an average score of 61.36, representing a relative improvement of +3.40 points over GRPO. Notably, the gains are especially pronounced on high-difficulty datasets such as AIME24 and AMC, where exploration plays a critical role in finding correct reasoning paths.
Compared to entropy-regularized variants, AEPO also exhibits clear advantages. Both Entropy-Reg and Entropy-Adv deliver only marginal improvements over GRPO. As shown in Fig. 4, the entropy regularization term in Entropy-Reg introduces significant bias: during mid-stage training, entropy begins to surge uncontrollably, indicating that entropy gradually replaces accuracy as the dominant optimization signal, with accumulated bias distorting the learning process. Entropy-Adv, on the other hand, fails to fundamentally reverse the trend of entropy collapse, leaving the model with insufficient exploration. In contrast, AEPO achieves consistent improvements across all benchmarks, underscoring the benefits of principled entropy control over heuristic entropy bonuses.
Figure 4: Entropy trajectories of AEPO compared with GRPO, Entropy-Reg, and Entropy-Adv. While GRPO and Entropy-Adv suffer from monotonic entropy collapse and fail to maintain exploration, Entropy-Reg leads to unstable fluctuations. In contrast, AEPO stabilizes entropy around the target level, demonstrating controllable and robust entropy regulation throughout training.
5.3 Entropy and exploration
Table 3: Comparison of different ablation losses: entropy control ability, and benchmark performance.
Loss and performance Entropy control
$\begin{aligned} \mathcal{J}(\theta)&=\;\mathcal{J}_{GRPO}(\theta)+\alpha\;\mathbb{E}_{q\sim P(Q),\;\{o_{i}\}_{i=1}^{G}\sim{\color[rgb]{0.828125,0.18359375,0.18359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.828125,0.18359375,0.18359375}\pi_{\theta_{\text{old}}}(O|q)}}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\Big[r_{i,t}(\theta)\,R(q,o_{i}),\;\text{clip}\!\big(r_{i,t}(\theta),1-\epsilon,1+\epsilon\big)\,R(q,o_{i})\Big]\end{aligned}$
AIME24 AMC Challenge Math GSM8K MATH Minerva Math Olympiad Avg
33.3 75.0 47.7 88.9 80.0 33.5 40.1 56.93 (-4.43)
entropy collapse
$\begin{aligned} \mathcal{J}(\theta)&=\;\mathcal{J}_{GRPO}(\theta)+\alpha\;\mathbb{E}_{q\sim P(Q),\;\{o_{i}\}_{i=1}^{G}\sim\pi^{T}_{\theta_{\text{old}}}(O|q)}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\Big[r_{i,t}(\theta)\,{\color[rgb]{0.828125,0.18359375,0.18359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.828125,0.18359375,0.18359375}\hat{A}_{i,t}},\;\text{clip}\!\big(r_{i,t}(\theta),1-\epsilon,1+\epsilon\big)\,{\color[rgb]{0.828125,0.18359375,0.18359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.828125,0.18359375,0.18359375}\hat{A}_{i,t}}\Big]\end{aligned}$
AIME24 AMC Challenge Math GSM8K MATH Minerva Math Olympiad Avg
36.7 75.0 48.3 88.1 80.8 36.0 41.8 58.14 (-3.22)
entropy collapse
To further examine the relationship between entropy and exploration, we evaluate AEPO under different entropy thresholds $\mathcal{H}$. As shown in Table 2, AEPO consistently surpasses GRPO across all benchmarks regardless of the entropy target, with strong improvements on challenging datasets such as AIME24 and MATH. At different thresholds, AEPO maintains clear advantages, indicating that principled entropy regulation yields benefits across a wide range of entropy levels.
Figure 1 further highlights this effect. While GRPO exhibits monotonic entropy collapse, AEPO stably maintains entropy around arbitrary preset levels simply by adjusting the hyperparameter $\mathcal{H}$, since AEPO is not sensitive to the parameter $\alpha$, as long as it is chosen within a reasonable range. This controllability allows us to probe the direct impact of entropy on exploration and downstream performance. Notably, we observe a non-monotonic trend: performance first improves as entropy rises from low to moderate levels, but begins to decline when entropy becomes excessively high. This provides explicit evidence that moderate entropy fosters effective exploration, while overly high entropy disperses optimization and harms accuracy.
Together, these results confirm that AEPO not only eliminates entropy collapse but also clarifies the link between entropy, exploration, and performance, offering a controllable pathway to balance exploration with convergence.
6 Ablation Study
To assess the contribution of each component in AEPO, we design two ablation studies to verify the necessity of distribution as regularization and REINFORCE as regularization for achieving effective entropy control.
6.1 Distribution as regularization
One critical component of AEPO is the use of temperature-adjusted distributions for entropy control. In our design, the REINFORCE regularization term samples from a modified distribution, where the temperature $T$ is adaptively adjusted based on the previous step’s entropy. This adjustment ensures that positive samples carry either higher or lower entropy as required, thereby stabilizing the overall entropy around the target threshold $\mathcal{H}$.
To validate the necessity of this design, we replace the temperature-adjusted distribution with the original distribution, as shown in Table 3. The results demonstrate that when REINFORCE samples are drawn directly from the original policy distribution (i.e., without temperature adjustment), entropy control collapses: the policy entropy monotonically decreases during training, similar to GRPO. More importantly, the average benchmark score drops to 57.39, which is even worse than standard GRPO and far below AEPO (61.36). This shows that the variance of REINFORCE gradients, when unregularized by distribution adjustment, further degrades optimization stability and downstream reasoning accuracy.
These findings confirm that distribution adjustment is indispensable for AEPO: it directly enables controllable entropy regulation, reduces variance, and consistently improves reasoning performance across benchmarks. Moreover, it provides strong evidence for the role of policy gradient as regularization: the REINFORCE term in AEPO influences the exploration ability of the GRPO component through entropy control, thereby shaping the overall optimization dynamics during training.
6.2 REINFORCE as regularization
Another essential component of AEPO is the use of REINFORCE gradients as a replacement for conventional entropy bonuses. In principle, REINFORCE allows unbiased estimation of gradients from positive samples while discarding negative ones, thereby forming a unidirectional optimization signal toward better distributions. This mechanism is critical for maintaining stable entropy control: when negative samples are included, the entropy-regularizing effect contributed by positive samples is counteracted, and the policy entropy eventually collapses.
To examine the role of REINFORCE, we conduct an ablation where the regularization term is still sampled from temperature-adjusted distributions, but the advantage function $\hat{A}_{t}$ is used directly without filtering negative samples. As shown in Table 3, this variant fails to prevent entropy collapse, leading to degraded performance across benchmarks. The average score drops to 58.14. This confirms that filtering negative samples via REINFORCE is indispensable.
7 Conclusion and discussion
In this paper, we propose Arbitrary Entropy Policy Optimization (AEPO), a principled RFT framework that addresses one of the most persistent challenges in RFT—precise and stable entropy control. Unlike traditional entropy regularization methods that trade off exploration against stability, AEPO achieves controllable entropy regulation through a unified design that integrates policy gradient, distribution, and REINFORCE as regularization components. This formulation eliminates the entropy collapse phenomenon in GRPO and maintains policy entropy within an arbitrarily specified range, enabling balanced and consistent exploration throughout training.
Extensive experiments across seven mathematical reasoning benchmarks demonstrate that AEPO consistently outperforms entropy-based baselines, exhibiting greater stability, generalization, and robustness to hyperparameters. More importantly, AEPO reveals a non-monotonic relationship between entropy and reasoning performance, showing that moderate entropy fosters exploration while excessive entropy impairs optimization—offering the first quantitative evidence linking entropy dynamics to reasoning capability in large language models.
Beyond solving entropy collapse, AEPO establishes a generalizable paradigm for learning under target distributions, suggesting its applicability to broader domains such as multi-modal alignment and long-horizon reasoning. In summary, AEPO transforms entropy from a passive statistical indicator into an active, tunable variable that fundamentally governs exploration, stability, and performance in RFT.
Declaration of AI
AI is only used for translation and language polishing in this paper.

### v4

@@
-Arbitrary Entropy Policy Optimization: Entropy Is Controllable in Reinforcement Finetuning
+Arbitrary Entropy Policy Optimization: Entropy Is Controllable in Reinforcement Fine-tuning
 Abstract
-Reinforcement finetuning (RFT) is essential for enhancing the reasoning capabilities of large language models (LLM), yet the widely adopted Group Relative Policy Optimization (GRPO) suffers from entropy collapse, where entropy monotonically decreases, exploration vanishes, and policies converge prematurely.
+Reinforcement fine-tuning (RFT) is essential for enhancing the reasoning capabilities of large language models (LLM), yet the widely adopted Group Relative Policy Optimization (GRPO) suffers from entropy collapse, where entropy monotonically decreases, exploration vanishes, and policies converge prematurely.
 Existing entropy-regularized methods only partially alleviate this issue while introducing bias and instability, leaving entropy control unresolved and the connection between entropy, exploration, and performance unclear.
@@
 Experiments demonstrate three major contributions: AEPO (1) stabilizes entropy at arbitrary target levels, effectively removing collapse in GRPO; (2) reveals a non-monotonic relation where performance first improves then declines with increasing entropy, clarifying the link between entropy, exploration, and reasoning; and (3) generalizes beyond entropy, providing a broader RFT paradigm where superior target distributions can serve as REINFORCE regularizers.
+Our code is available at https://github.com/597358816/AEPO.
 Machine Learning, ICML
@@
 1 Introduction
-Reinforcement finetuning (RFT) has become a cornerstone for enhancing the reasoning capabilities of large language models (LLM) (GLM et al., 2024; Touvron et al., 2023; Schulman et al., 2017; Rafailov et al., 2023; Zhong et al., 2024; Wang et al., 2024).
+Reinforcement fine-tuning (RFT) has become a cornerstone for enhancing the reasoning capabilities of large language models (LLM) (GLM et al., 2024; Touvron et al., 2023; Schulman et al., 2017; Rafailov et al., 2023; Zhong et al., 2024; Wang et al., 2024).
 Among existing approaches, Group Relative Policy Optimization (GRPO) has gained wide adoption due to its efficiency and scalability (Shao et al., 2024; Liu et al., 2024; Guo et al., 2025).
@@
 AEPO consistently outperforms all baselines, achieving stable entropy control and delivering substantial improvements across diverse benchmarks, with the best overall performance at $\mathcal{H}=0.75$.
-Benchmarks AIME24 AMC Challenge math GSM8K MATH Minerva math Olympiad Average
+Benchmarks AIME24 AMC College Math GSM8K MATH Minerva Math Olympiad Average
 Qwen2.5-math-7B 13.3 40.0 41.7 65.4 65.5 11.0 26.7 37.66
@@
 For training, we use the DAPO-17K dataset (Yu et al., 2025), which contains diverse problem instances curated for RFT.
-Benchmark: Evaluation is performed on a broad suite of mathematical reasoning benchmarks, including AIME24 (HuggingFaceH4, 2025), AMC (Lightman et al., 2023), College Math (Zhong et al., 2023), GSM8K (Cobbe et al., 2021), MATH (Lightman et al., 2023), Minerva Math (Lewkowycz et al., 2022), and Olympiad (Lightman et al., 2023).
+Benchmark: Evaluation is performed on a broad suite of mathematical reasoning benchmarks, including AIME24 (HuggingFaceH4, 2025), AMC (Lightman et al., 2023), College Math (Tang et al., 2024), GSM8K (Cobbe et al., 2021), MATH (Lightman et al., 2023), Minerva Math (Lewkowycz et al., 2022), and Olympiad (Lightman et al., 2023).
 These benchmarks collectively span a wide range of difficulty levels, from grade school arithmetic to advanced competition-level mathematics, and together they cover nearly all mainstream benchmarks for mathematical reasoning, enabling a comprehensive assessment of AEPO’s impact on reasoning performance across diverse tasks.
@@
 $\begin{aligned} \mathcal{J}(\theta)&=\;\mathcal{J}_{GRPO}(\theta)+\alpha\;\mathbb{E}_{q\sim P(Q),\;\{o_{i}\}_{i=1}^{G}\sim{\color[rgb]{0.828125,0.18359375,0.18359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.828125,0.18359375,0.18359375}\pi_{\theta_{\text{old}}}(O|q)}}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\Big[r_{i,t}(\theta)\,R(q,o_{i}),\;\text{clip}\!\big(r_{i,t}(\theta),1-\epsilon,1+\epsilon\big)\,R(q,o_{i})\Big]\end{aligned}$
-AIME24 AMC Challenge Math GSM8K MATH Minerva Math Olympiad Avg
+AIME24 AMC College GSM8K MATH Minerva Math Olympiad Avg
 33.3 75.0 47.7 88.9 80.0 33.5 40.1 56.93 (-4.43)
@@
 $\begin{aligned} \mathcal{J}(\theta)&=\;\mathcal{J}_{GRPO}(\theta)+\alpha\;\mathbb{E}_{q\sim P(Q),\;\{o_{i}\}_{i=1}^{G}\sim\pi^{T}_{\theta_{\text{old}}}(O|q)}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\Big[r_{i,t}(\theta)\,{\color[rgb]{0.828125,0.18359375,0.18359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.828125,0.18359375,0.18359375}\hat{A}_{i,t}},\;\text{clip}\!\big(r_{i,t}(\theta),1-\epsilon,1+\epsilon\big)\,{\color[rgb]{0.828125,0.18359375,0.18359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.828125,0.18359375,0.18359375}\hat{A}_{i,t}}\Big]\end{aligned}$
-AIME24 AMC Challenge Math GSM8K MATH Minerva Math Olympiad Avg
+AIME24 AMC College GSM8K MATH Minerva Math Olympiad Avg
 36.7 75.0 48.3 88.1 80.8 36.0 41.8 58.14 (-3.22)

### v6

@@
-Arbitrary Entropy Policy Optimization: Entropy Is Controllable in Reinforcement Fine-tuning
+Arbitrary Entropy Policy Optimization Breaks The Exploration Bottleneck of Reinforcement Learning
 Abstract
-Reinforcement fine-tuning (RFT) is essential for enhancing the reasoning capabilities of large language models (LLM), yet the widely adopted Group Relative Policy Optimization (GRPO) suffers from entropy collapse, where entropy monotonically decreases, exploration vanishes, and policies converge prematurely.
-Existing entropy-regularized methods only partially alleviate this issue while introducing bias and instability, leaving entropy control unresolved and the connection between entropy, exploration, and performance unclear.
-We propose Arbitrary Entropy Policy Optimization (AEPO), which eliminates entropy collapse by replacing entropy bonuses with REINFORCE policy gradient on temperature-adjusted distributions and stabilizing entropy through temperature regulation.
-AEPO integrates three key designs: policy gradient as regularization, distribution as regularization, and REINFORCE as regularization, enabling precise entropy control without distorting optimization.
-Experiments demonstrate three major contributions: AEPO (1) stabilizes entropy at arbitrary target levels, effectively removing collapse in GRPO; (2) reveals a non-monotonic relation where performance first improves then declines with increasing entropy, clarifying the link between entropy, exploration, and reasoning; and (3) generalizes beyond entropy, providing a broader RFT paradigm where superior target distributions can serve as REINFORCE regularizers.
-Our code is available at https://github.com/597358816/AEPO.
+Reinforcement Learning (RL) is essential for enhancing the reasoning capabilities of large language models (LLMs), yet the widely adopted Group Relative Policy Optimization (GRPO) suffers from entropy collapse, causing exploration to vanish and policies to converge prematurely.
+As a result, RL is widely believed to be incapable of expanding the reasoning frontier of LLMs.
+Existing entropy-regularized methods introduce an inevitable trade-off between reward and entropy, leading to exploration accompanied by non-negligible optimization bias.
+In this work, we prove that temperature-guided REINFORCE can modulate policy entropy, and propose Arbitrary Entropy Policy Optimization (AEPO), which reformulates entropy regularization as a policy-gradient optimization problem.
+Rather than manipulating entropy directly, AEPO implicitly regulates it by applying a REINFORCE regularization term on temperature-adjusted samples, ensuring that entropy is controlled but never dominates optimization, thereby enabling arbitrary and principled entropy regulation.
+Experiments show that AEPO outperforms RL baselines on both pass@1 and pass@$k$, and even surpasses the base model on pass@1024.
+By modulating entropy precisely, AEPO achieves more effective optimization dynamics and provides direct empirical evidence that entropy, exploration, and performance are intrinsically linked.
 Machine Learning, ICML
@@
 inkscapearea=page
-Figure 1: Entropy across five runs of AEPO. By adjusting only the parameter $\mathcal{H}$, entropy can be stably maintained at different levels.
+Figure 1: Entropy across five runs of AEPO. By adjusting only the parameter $\mathcal{H}$, entropy can be controlled at different levels.
 1 Introduction
-Reinforcement fine-tuning (RFT) has become a cornerstone for enhancing the reasoning capabilities of large language models (LLM) (GLM et al., 2024; Touvron et al., 2023; Schulman et al., 2017; Rafailov et al., 2023; Zhong et al., 2024; Wang et al., 2024).
-Among existing approaches, Group Relative Policy Optimization (GRPO) has gained wide adoption due to its efficiency and scalability (Shao et al., 2024; Liu et al., 2024; Guo et al., 2025).
-However, GRPO suffers from a well-documented drawback: entropy collapse: As training progresses, policy entropy declines monotonically, sampled outputs converge to nearly identical solutions, and the model prematurely adopts a deterministic policy with limited exploration (Yu et al., 2025; Li et al., 2025a; Zhang et al., 2025).
-This severely restricts the ability of RFT to discover diverse reasoning strategies.
-Existing remedies typically rely on entropy regularization and clip refine, which inevitably introduce bias and instability into the optimization objective, creating a trade-off between exploration and stability (Hou et al., 2025; Cui et al., 2025; Cheng et al., 2025; Shen, 2025).
-EFRame is the first to achieve large-scale entropy control, but it remains a qualitative analysis without conducting a quantitative study of entropy (Wang et al., 2025a).
-Although the issue of entropy collapse in GRPO has been repeatedly noted, it remains unresolved: there is still no principled algorithm capable of precisely regulating entropy throughout training.
-Existing approaches often treat entropy only as a side indicator—for instance, using it to split steps or guide auxiliary heuristics—while the policy entropy itself continues to collapse as training progresses (Wang et al., 2025b; Zheng et al., 2025; Li et al., 2025b; Cheng et al., 2025).
-As a result, the relationship among entropy, exploration, and performance remains unclear: it is uncertain whether entropy is a sufficient proxy for exploration or whether exploration itself consistently improves training outcomes.
+Reinforcement Learning (RL) has become a cornerstone for enhancing the reasoning capabilities of large language models (LLM) (glm2024chat; touvron2023llama; schulman2017proximal; rafailov2023direct; zhong2024dpo; wang2024comprehensive).
+Entropy reflects the extent of exploration by measuring the model’s uncertainty and output diversity during reasoning (schulman2017equivalence; haarnoja2018soft; nachum2017bridging).
+Among existing approaches, Group Relative Policy Optimization (GRPO) has gained wide adoption due to its efficiency and scalability (shao2024deepseekmath; liu2024deepseek; guo2025deepseek).
+However, GRPO suffers from a well-documented drawback: entropy collapse: As training progresses, policy entropy declines monotonically, sampled outputs converge to nearly identical solutions, and the model prematurely adopts a deterministic policy with limited exploration (yu2025dapo; li2025disco; zhang2025edge).
+This limits the ability of RL to discover diverse reasoning strategies.
+Recent studies have shown that RL fails to broaden the reasoning capabilities of LLMs; instead, it merely sharpens behaviors within the base model’s existing knowledge without introducing genuinely new reasoning patterns (yue2025does).
+This highlights the necessity of effective exploration in RL for LLMs, and such exploration fundamentally requires breaking the entropy bottleneck (cui2025entropy).
+Although the problem of entropy collapse in GRPO has been repeatedly recognized, it remains fundamentally unsolved.
+Most existing variants incorporate entropy as an additional reward term in the gradient, such as entropy regularization (hou2025advancing; cui2025entropy; shen2025entropy) or advantage-weighted bonuses (cheng2025reasoning), to partially alleviate entropy collapse.
+However, these methods introduce an inevitable trade-off between reward and entropy, and thus bring non-negligible optimization bias when promoting exploration.
+In practice, entropy may oscillate between collapse and explosion instead of stabilizing in an optimal exploration regime, making the entropy–exploration–performance relationship hard to observe.
+As a result, it remains unclear whether entropy is a sufficient proxy for exploration and whether exploration itself consistently improves training outcomes.
 If significant performance improvement can be observed by adjusting entropy to a better range, it would indicate that exploration plays a crucial role in this process.
-If arbitrary entropy control can be achieved during RFT, it would make it possible to realize exploration at any desired degree and, in turn, establish a principled connection among entropy, exploration, and performance.
-Motivated by these challenges, we propose Arbitrary Entropy Policy Optimization (AEPO), a novel policy gradient that directly addresses entropy collapse.
-AEPO replaces conventional entropy bonuses with a REINFORCE policy gradient (Williams, 1992; Sutton et al., 1999) applied to samples drawn from temperature-adjusted distributions.
-Entropy is further stabilized through temperature regulation, ensuring that batch-level entropy remains oscillating around an arbitrary constant $\mathcal{H}$ throughout training, as shown in Fig.
-1.
+If arbitrary entropy control can be achieved during RL, it would make it possible to realize exploration at any desired degree and, in turn, establish a principled connection among entropy, exploration, and performance.
+Motivated by these challenges, we propose Arbitrary Entropy Policy Optimization (AEPO), which reformulates entropy regularization as a policy optimization, thereby fundamentally resolving entropy collapse.
+Instead of adding entropy bonuses, AEPO applies a REINFORCE policy gradient to temperature-adjusted samples, avoiding the unestimable bias introduced by conventional entropy-regularized methods.
+As shown in Fig.
+1, AEPO maintains entropy oscillation around an arbitrary constant and is even theoretically capable of approximating arbitrary functions.
 AEPO achieves entropy control through three key design components:
-  • Policy gradient as regularization: Instead of relying on a conventional entropy bonus, we employ a full policy gradient term applied to samples that naturally exhibit high- or low-entropy properties.
-This design ensures that entropy never dominates the optimization objective, allowing the model to monotonically explore toward higher accuracy throughout training.
-  • Distribution as regularization: Entropy is regulated via temperature-adjusted distributions.
-Given a predefined entropy threshold $\mathcal{H}$, when the previous step’s entropy $\mathcal{H}(\pi_{\theta_{\text{old}}})<\mathcal{H}$, we regard the distribution under $T_{\text{high}}>1$ as a better candidate and mix in a small proportion of its samples.
-Conversely, when $\mathcal{H}(\pi_{\theta_{\text{old}}})\geq\mathcal{H}$, we instead sample from the distribution under $T_{\text{low}}<1$.
-  • REINFORCE as regularization: In Reinforcement Learning with Verifiable Reward (RLVR), the REINFORCE algorithm can filter out negative samples in an unbiased manner, allowing positive samples to form a unidirectional gradient toward a better distribution.
-This guides the model to optimize from positive samples that align with the target distribution.
-If negative samples were included at this stage, the entropy control mechanism would fail.
+  • Policy gradient as regularization: Instead of using an entropy bonus, AEPO applies a full policy-gradient term to samples with naturally high or low entropy, preventing entropy from dominating optimization while enabling stable exploration.
+  • Temperature as regularization: Entropy is modulated through temperature-based sampling.
+When the current entropy falls below the target level, AEPO draws higher-temperature samples to increase entropy; when it exceeds the target, AEPO instead draws lower-temperature samples to reduce entropy.
+  • REINFORCE as regularization: In RLVR, the REINFORCE algorithm can filter out negative samples without introducing bias, allowing positive samples to form a unidirectional gradient toward a better distribution.
 In summary, our contributions are threefold:
-  • Controllable entropy: We propose Arbitrary Entropy Policy Optimization that can stabilize entropy at arbitrary target levels, effectively eliminating entropy collapse in GRPO.
+  • Breaking the entropy bottleneck: We propose Arbitrary Entropy Policy Optimization, which can stabilize entropy at arbitrary target levels, effectively eliminating entropy collapse in GRPO and enabling exploration beyond the longstanding entropy bottleneck.
+AEPO introduces a REINFORCE regularization mechanism that controls entropy without distorting the optimization objective, representing a fundamentally new paradigm for exploration.
+  • Breaking exploration bottleneck: AEPO achieves consistent improvements over RL baselines on both pass@1 and pass@k, and even surpasses the base model on pass@1024, demonstrating that RL can conduct exploration rather than merely sharpening the base model’s knowledge.
   • Entropy–performance relation: We find that merely adjusting entropy can directly influence training performance, providing explicit evidence for the correlation between entropy, exploration, and performance.
 Moreover, we observe a non-monotonic trend in which performance first increases and then decreases as entropy grows, highlighting the existence of an optimal entropy regime.
-  • Generalizability: Beyond entropy control, AEPO provides a broader paradigm for RFT.
-When a target distribution $\pi^{*}$ is identified as superior to the current policy $\pi_{\theta}$, samples from $\pi^{*}$ can be used to construct a REINFORCE policy gradient as a regularization term, which allows $\pi$ to progressively approximate $\pi^{*}$ during long-horizon training.
 2 Related Work
-RFT has become a central paradigm for post-training large language models (LLMs), aligning them with human feedback and task-specific objectives.
-Early methods such as RLHF leveraged policy optimization (e.g., PPO) to encode human preferences (OpenAI, 2023; Team et al., 2024; Wei et al., 2023; Liu et al., 2023), while Direct Preference Optimization (DPO) (Rafailov et al., 2024) later improved efficiency by optimizing policies directly from preference data.
-Recent models like DeepSeek-R1 (Guo et al., 2025) and Kimi-1.5 (Team et al., 2025) extend RFT through hybrid reward formulations and scalable optimization.
-Among them, Group Relative Policy Optimization (GRPO) (Shao et al., 2024; Liu et al., 2024) has become the de facto baseline for reasoning-focused RFT, yet suffers from entropy collapse that limits exploration of diverse reasoning strategies.
+RL has become a central paradigm for post-training large language models (LLMs), aligning them with human feedback and task-specific objectives.
+Early methods, such as RLHF, leveraged policy optimization (e.g., PPO) to encode human preferences (openai2023gpt4; team2024gemini1_5; wei2023instructiongpt; liu2023visual), while Direct Preference Optimization (DPO) (rafailov2024direct) later improved efficiency by optimizing policies directly from preference data.
+Recent models like DeepSeek-R1 (guo2025deepseek), and Kimi-1.5 (team2025kimi1_5) extend RL through hybrid reward formulations and scalable optimization.
+Among them, Group Relative Policy Optimization (GRPO) (shao2024deepseekmath; liu2024deepseek) has become the de facto baseline for reasoning-focused RL, yet suffers from entropy collapse that limits exploration of diverse reasoning strategies.
 Entropy has long been regarded as a proxy for exploration in reinforcement optimization.
-Classical methods employ entropy regularization to stabilize training and encourage diversity (Sutton et al., 1999; Williams, 1992).
-More recent studies extend this idea to GRPO by introducing entropy bonuses into rewards or advantages (Cheng et al., 2025; Cui et al., 2025; Shen, 2025).
-However, such approaches only yield coarse-grained effects: entropy still collapses as training proceeds, or the added bias destabilizes optimization.
-EFRame (Wang et al., 2025a) represents the first framework to achieve large-scale entropy control through conditional sampling and replay, yet its treatment remains qualitative and lacks quantitative characterization of the entropy–exploration–performance relationship.
+Classical methods employ entropy regularization to stabilize training and encourage diversity (sutton1999policy; williams1992simple).
+More recent studies extend this idea to GRPO by introducing entropy bonuses into rewards or advantages (cheng2025reasoning; cui2025entropy; shen2025entropy).
+However, such approaches only yield coarse-grained effects: entropy still collapses as training proceeds, or the added bias distorts optimization.
 In summary, existing methods lack a principled mechanism to precisely regulate entropy throughout training.
@@
 3 Preliminary
-Our work focuses on fine-tuning LLM using Reinforcement Learning (RL) for tasks with verifiable solutions, such as mathematical reasoning and code generation.
+Our work focuses on fine-tuning LLM using Reinforcement Learning with Verifiable Reward (RLVR), such as mathematical reasoning and code generation.
 Suppose the LLM is a softmax policy, that is
@@
 $\mathbb{E}_{q\sim P(Q),o\sim\pi_{\theta}(O|q)}\sum_{t=1}^{|o|}[R(q,o)-\lambda\cdot\text{log}\pi_{\theta}(o_{t}|q,o_{<t})].$
-Entropy-Adv(Cheng et al., 2025) proposed an entropy-guided advantage shaping method.
+Entropy-Adv(cheng2025reasoning) proposed an entropy-guided advantage shaping method.
 The key idea is to inject an entropy-based term into the advantage function during policy optimization.
@@
 The policy gradient of the algorithm retains a format similar to that in GRPO, where only the advantage $\hat{A}_{i,t}$ is replaced by the shaped one:
-$\nabla_{\theta}\mathcal{J}^{\text{shaped}}(\theta)=$ $\mathbb{E}_{q\sim P(Q),o\sim\pi_{\theta}(O|q)},$
+$\nabla_{\theta}\mathcal{J}^{\text{shaped}}(\theta)=$ $\mathbb{E}_{q\sim P(Q),o\sim\pi_{\theta}(O|q)}$
 $\sum_{t=1}^{|o|}[\nabla_{\theta}\text{log}\pi_{\theta}(o_{t}|q,o_{<t})\cdot A_{i,t}^{\text{shaped}}].$
-Table 1: Comparison of entropy-based optimization objectives.
-Policy Gradient
-Entropy-Reg $\sum_{t}(\hat{A}_{t}\cdot\nabla_{\theta}\text{log}\pi_{\theta}(o_{t}|q,o_{<t})+\lambda\cdot\nabla_{\theta}\mathcal{H}_{t})$
-Entropy-Adv $\sum_{t}A_{i,t}^{\text{shaped}}\cdot\nabla_{\theta}\text{log}\pi_{\theta}(o_{t}|q,o_{<t})$
 4 Method
-Arbitrary Entropy Policy Optimization (AEPO) is designed to achieve precise and stable control of policy entropy during RFT.
-Unlike conventional entropy-regularized methods, AEPO does not introduce explicit entropy bonuses; instead, it adjusts the training distribution and policy gradient to regulate entropy implicitly yet controllably.
-This section first introduces the key premises underlying AEPO’s design, which reveal the relationship between temperature, entropy, and policy updates, and then details the algorithmic formulation of AEPO.
-$\mathcal{J}_{GRPO}(\theta)=\;\mathbb{E}_{q\sim P(Q),\;\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O|q)}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\Big[r_{i,t}(\theta)\,\hat{A}_{i,t},\;\text{clip}\!\big(r_{i,t}(\theta),1-\epsilon,1+\epsilon\big)\,\hat{A}_{i,t}\Big],$ (1)
-$\mathcal{J}_{AEPO}(\theta)=\;\mathcal{J}_{GRPO}(\theta)+\alpha\;\mathbb{E}_{q\sim P(Q),\;\{o_{i}\}_{i=1}^{G}\sim\pi^{T}_{\theta_{\text{old}}}(O|q)}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\Big[r_{i,t}(\theta)\,R(q,o_{i}),\;\text{clip}\!\big(r_{i,t}(\theta),1-\epsilon,1+\epsilon\big)\,R(q,o_{i})\Big],$
+Arbitrary Entropy Policy Optimization (AEPO) is designed to achieve precise and stable control of policy entropy during RL.
+This section first introduces the key theoretical premises underlying AEPO’s design, which reveal the relationship between temperature, entropy, and policy updates, and then details the algorithmic formulation of AEPO.
+$\begin{split}\mathcal{J}_{GRPO}(\theta)=&\;\mathbb{E}_{q\sim P(Q),\;\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O|q)}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\Big[r_{i,t}(\theta)\,\hat{A}_{i,t},\;\text{clip}\!\big(r_{i,t}(\theta),1-\epsilon,1+\epsilon\big)\,\hat{A}_{i,t}\Big],\\[5.69054pt] \mathcal{J}_{AEPO}(\theta)=&\mathcal{J}_{GRPO}+\alpha\;\mathbb{E}_{q\sim P(Q),\;\{o_{i}\}_{i=1}^{G}\sim\pi^{T}_{\theta_{\text{old}}}(O|q)}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\Big[r_{i,t}(\theta)\,R(q,o_{i}),\;\text{clip}\!\big(r_{i,t}(\theta),1-\epsilon,1+\epsilon\big)\,R(q,o_{i})\Big],\end{split}$ (1)
 where $r_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}\mid q)}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid q)}$ and $T=T_{\text{low}}+\big(T_{\text{high}}-T_{\text{low}}\big)\,\mathbf{1}\!\left[\,\mathcal{H}(\pi_{\theta_{\text{old}}})<\mathcal{H}\,\right]$.
-4.1 Premise
+4.1 Theoretical analysis
 The design of AEPO is built upon two empirical premises that connect temperature-based sampling to entropy dynamics.
-These premises establish the foundation for controllable entropy modulation without introducing bias into the optimization objective.
-Premise 1.
+These analyses establish the foundation for controllable entropy modulation without introducing bias into the optimization objective.
+Lemma 4.1.
 Higher temperature distributions globally correspond to higher policy entropy, while lower temperature corresponds to lower entropy.
-Previous studies, such as Du et al.
-(2025), show that increasing the sampling temperature generally broadens the model’s output distribution and raises its entropy.
-Although the correspondence between higher temperature and higher entropy is not a universally proven principle, it holds—or approximately holds—under most intuitive and commonly used model settings.
-In the context of RFT, temperature thus serves as a practical external variable for indirectly regulating entropy, enabling controllable entropy adjustment.
-Figure 2: Entropy dynamics under temperature-controlled sampling.
-High-temperature positive samples increase entropy, promoting exploration, while low-temperature positive samples reduce entropy, leading to more deterministic behavior.
-Lemma 4.1.
-(Cui et al., 2025) If the actor policy $\pi_{\theta}$ is a tabular softmax policy updated via natural policy gradient (Kakade, 2001) with step size $\eta$.
-Then the change in policy entropy between two steps approximately satisfies:
-$\begin{split}\mathcal{H}(\pi_{\theta}^{k+1})-\mathcal{H}(\pi_{\theta}^{k})\approx&-\eta\cdot\mathbb{E}_{s\sim d_{\mu}^{k}}\text{Cov}_{a\sim\pi^{k}_{\theta}(\cdot\mid s)}\\ &\!\bigl[\log\!\pi^{k}_{\theta}(a\mid s),\,A^{\pi^{k}}(s,a)\bigr].\end{split}$
-The proof can be seen in Liu (2025) and Cui et al.
-(2025).
-$\mathcal{H}$ indicates the policy entropy of policy model, and Cov denotes covariance, $\pi^{k}_{\theta}$ is the policy at step $k$, and $A^{\pi^{k}}(s,a)$ is the advantage function of action $a$ under state $s$.
-This result indicates that when positive actions have a low probability.
-If the probability of sampling low-probability actions can be increased, the model will be optimized toward higher entropy.
-Premise 2.
-Positive samples from temperature-adjusted distributions induce predictable entropy change during training.
-When positive samples are drawn from a higher-temperature version of the current policy distribution, training with these samples increases the model’s entropy, promoting exploration.
-Conversely, sampling from a lower-temperature distribution produces entropy reduction, guiding the model toward more deterministic behavior.
-Figure 2 illustrates this phenomenon: the entropy increases under high-temperature sampling and decreases under low-temperature sampling.
-Meanwhile, entropy increase is typically slower than entropy decrease, which aligns with intuition.
-This dynamic forms the empirical basis for AEPO’s temperature-controlled entropy feedback loop, enabling bidirectional regulation of entropy around a target value.
+Previous studies, such as du2025optimizing, show that increasing the sampling temperature generally broadens the model’s output distribution and raises its entropy.
+In the context of RL, temperature can adjust entropy during inference, but it does not modify the underlying policy itself.
+Simply increasing temperature to force diverse outputs leads to off-policy sampling, which undermines the consistency required for policy optimization.
+What is needed instead is a mechanism that directly adjusts the policy itself so that the original policy is capable of generating more diverse behaviors.
+Assumption 4.2.
+Assume that $T>1$ and the actor policy $\pi_{\theta}$ is a tabular softmax policy updated via the following equation:
+$\theta_{k+1}=\theta_{k}+\eta\mathbb{E}_{a\sim\pi_{\theta_{k}}^{T}(\cdot|s)}[R(s,a)\cdot\nabla_{\theta}\text{log}\pi_{\theta_{k}}(a|s)].$
+Theorem 4.3.
+Denote the function
+$\Delta H_{k}(T)=-\eta\cdot$ $\sum_{a^{*}\in\mathcal{A}^{*}}\pi_{\theta_{k}}^{T}(a^{*}|s)\times$
+$\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a^{*}}-\pi_{\theta_{k}}(a^{\prime}|s)).$
+Then we have
+$\Delta H_{k}(T)\approx H(\theta_{k+1})-H(\theta_{k}),$
+and
+$\Delta H_{k}^{{}^{\prime}}(T)|_{T=1}>0\quad\text{if}\quad\Delta H_{k}(T)|_{T=1}<0.$
+The proof is provided in Appendix B.5.
+In more intuitive terms, the preceding theorem implies that if the policy entropy begins to collapse during training, we can counteract this downward trend by performing updates using REINFORCE policy gradient of high-temperature samples.
+Corollary 4.4.
+High-temperature REINFORCE induces a relative increase in policy entropy, while low-temperature REINFORCE induces a relative decrease in policy entropy.
+Within the RLVR framework, the binary nature of the reward signal gives rise to an inherent filtering mechanism for negative samples.
+Building upon this property and the preceding premises, we can formulate a clear insight: computing the policy gradient with temperature-adjusted samples enables predictable and controllable entropy modulation.
+As illustrated in Figure 2, entropy increases under high-temperature sampling and decreases under low-temperature sampling, with the increase typically occurring more gradually than the decrease.
+This dynamic forms the empirical foundation for AEPO’s temperature-controlled entropy feedback loop, allowing bidirectional regulation of entropy around a target value.
+Figure 2: Entropy dynamics under temperature-controlled REINFORCE. High-temperature REINFORCE increases entropy, promoting exploration, while low-temperature REINFORCE reduces entropy.
 4.2 AEPO
-Building upon the premises introduced above, AEPO is designed to achieve stable and controllable entropy regulation throughout RFT.
-Its core idea is to replace explicit entropy bonuses with a REINFORCE regularization term that adjusts the sampling distribution according to the current entropy state.
-In this way, AEPO modulates exploration implicitly—through data distribution—rather than directly through the loss function.
-Formally, AEPO augments the standard GRPO objective with an additional policy gradient term applied to samples drawn from temperature-adjusted distributions.
-The overall policy gradient can be expressed as:
+Building on the theoretical premises described above, AEPO achieves stable and controllable entropy regulation by replacing explicit entropy bonuses with a REINFORCE-based regularization mechanism that adjusts the sampling temperature according to the current entropy state.
+Concretely, AEPO augments the GRPO objective (without KL divergence) with an additional policy gradient term applied to temperature-adjusted samples.
+The resulting policy gradient is given by:
 $\nabla_{\theta}\mathcal{J}_{AEPO}(\theta)$
@@
 As illustrated in Fig.
-3, entropy regularization alone tends to drive the optimization toward two extremes—either entropy collapse or entropy explosion.
+3, entropy regularization tends to drive the optimization toward two extremes—either entropy collapse or entropy explosion.
 In the former case, the regularization term is too weak to reverse the monotonic entropy decay; in the latter, entropy becomes an irreversible dominant factor in optimization.
@@
 Entropy regularization often drives optimization toward two extremes—collapse or explosion—while AEPO maintains entropy within a stable and optimal exploration range.
-Distribution as regularization.
-AEPO regulates the optimization direction by adjusting the expected sampling distribution $\pi^{*}$ derived from the current policy.
+Temperature as regularization.
+AEPO regulates the optimization direction by adjusting the sampling temperature.
 When the observed entropy $\mathcal{H}(\pi_{\theta_{\text{old}}})$ is below the target threshold $\mathcal{H}$, AEPO samples from the higher-temperature distribution $\pi^{T_{\text{high}}}_{\text{old}}$ to encourage exploration.
@@
 This property allows REINFORCE to filter out negative samples in an unbiased manner, ensuring that the gradient is formed from positive samples that align with the desired distribution.
-Consequently, AEPO’s regularization term produces a one-sided optimization signal that guides the policy toward higher-quality behavior distributions.
-Table 2: Main results on mathematical reasoning benchmarks.
-We compare Qwen2.5-math-7B (base model), GRPO, entropy-based baselines, and AEPO under different entropy thresholds $\mathcal{H}$.
-AEPO consistently outperforms all baselines, achieving stable entropy control and delivering substantial improvements across diverse benchmarks, with the best overall performance at $\mathcal{H}=0.75$.
-Benchmarks AIME24 AMC College Math GSM8K MATH Minerva Math Olympiad Average
-Qwen2.5-math-7B 13.3 40.0 41.7 65.4 65.5 11.0 26.7 37.66
-GRPO 36.7 75.0 48.2 88.9 80.5 34.6 41.8 57.96
-Entropy-Reg 36.7 75.0 47.6 87.0 80.4 35.7 40.4 57.39
-Entropy-Adv 36.7 75.0 47.8 87.8 80.4 37.5 42.1 58.18
-AEPO $\mathcal{H}=0.25$ 40.0 77.5 48.5 89.4 80.6 33.5 42.2 58.81
-AEPO $\mathcal{H}=0.50$ 43.3 82.5 48.8 89.5 81.6 38.2 43.0 60.99
-AEPO $\mathcal{H}=0.75$ 50.0 (+36.7) 80.0 (+40.0) 48.2 (+6.50) 89.4 (+24.0) 82.0 (+16.5) 37.5 (+26.5) 42.4 (+15.7) 61.36 (+23.70)
-AEPO $\mathcal{H}=1.00$ 33.3 75.0 48.4 88.7 80.8 37.9 42.1 58.03
+Consequently, AEPO’s regularization term produces an unidirectional optimization signal that guides the policy toward higher-quality behavior distributions.
+Figure 4: Entropy trajectories of AEPO compared with GRPO, Entropy-Reg, and Entropy-Adv.
+AEPO stabilizes entropy around a moderate level, demonstrating controllable and robust entropy regulation throughout training.
+Table 1: AEPO demonstrates consistently superior performance across all mathematical reasoning benchmarks, surpassing GRPO, and all entropy-regularized baselines.
+Benchmarks AIME24×32 AIME25×32 AMC×32 GSM8K MATH Minerva Olympiad Average
+Qwen2.5-7B 7.91 5.31 36.2 88.5 64.4 22.0 29.3 36.24
++GRPO 17.1 7.60 65.8 92.3 75.6 36.8 38.8 47.70
++Entropy-Reg 13.6 8.85 67.4 92.3 76.8 35.5 39.1 47.65
++Entropy-Adv 14.8 8.23 67.3 91.9 76.6 38.2 37.5 47.79
++AEPO    $\Delta$ vs. GRPO 17.5 (+0.4) 11.4 (+3.8) 69.3 (+3.5) 92.9 (+0.6) 78.0 (+2.4) 37.8 (+1.0) 40.2 (+1.4) 49.57 (+1.87)
+Qwen2.5-math-7B 15.5 7.81 42.1 65.4 59.4 11.0 26.7 32.56
++GRPO 32.1 11.0 72.4 88.7 80.6 34.6 41.8 51.60
++Entropy-Reg 31.4 10.1 74.3 87.0 78.8 35.7 40.4 51.10
++Entropy-Adv 32.1 11.4 72.1 87.8 78.8 37.5 42.1 51.76
++AEPO    $\Delta$ vs. GRPO 36.4 (+4.3) 12.6 (+1.6) 74.8 (+2.4) 89.5 (+0.8) 82.6 (+2.0) 38.2 (+3.6) 43.0 (+1.2) 53.87 (+2.27)
+Qwen3-4B 36.4 22.7 71.9 93.9 84.8 42.3 47.2 57.03
++GRPO 52.9 41.5 86.1 95.2 92.0 46.7 60.0 67.77
++AEPO    $\Delta$ vs. GRPO 54.5 (+1.6) 43.7 (+2.2) 89.7 (+3.6) 95.0 (-0.2) 92.8 (+0.8) 47.8 (+1.1) 60.9 (+0.9) 69.20 (+1.43)
+Table 2: Comparison of pass@$k$ performance between the base model, GRPO, and AEPO across four mathematical reasoning benchmarks (steps $\geq$ 100).
+The accompanying figures show the evolution of pass@1024 of AEPO and GRPO on AIME24 and AIME25 during training.
+Benchmarks AIME24 AIME25 Minerva Olympiad
+Pass@512 Pass@1024 Pass@512 Pass@1024 Pass@128 Pass@128
+Qwen2.5-7B 71.7 80.0 61.7 73.3 69.8 77.8
++GRPO 63.3 73.3 60.0 63.3 65.8 72.9
++AEPO 70.0 83.3 66.7 73.3 67.0 74.1
+Table 3: Ablation study on AEPO under different entropy targets, including benchmark performance across seven mathematical reasoning datasets.
+The upper table compares AEPO at various entropy levels, while the lower ablations illustrate how different loss formulations affect both performance and entropy stability.
+Benchmarks AIME24×32 AIME25×32 AMC×32 GSM8K MATH500 Minerva Olympiad Average
+Qwen2.5-math-7B 15.5 7.81 42.1 65.4 59.4 11.0 26.7 37.66
++AEPO $\mathcal{H}=0.25$ 37.9 11.3 74.4 89.4 79.8 36.0 42.2 53.00
++AEPO $\mathcal{H}=0.50$ 36.4 12.6 74.8 89.5 82.6 38.2 43.0 53.87
++AEPO $\mathcal{H}=0.75$ 33.6 15.1 77.0 89.4 79.2 37.5 42.4 53.45
++AEPO $\mathcal{H}=1.00$ 33.2 15.6 74.8 88.7 79.6 37.9 42.1 53.13
+Ablation loss and performance Entropy control
+$\begin{aligned} \mathcal{J}(\theta)&=\;\mathcal{J}_{GRPO}(\theta)+\alpha\;\mathbb{E}_{q\sim P(Q),\;\{o_{i}\}_{i=1}^{G}\sim{\color[rgb]{0.828125,0.18359375,0.18359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.828125,0.18359375,0.18359375}\pi_{\theta_{\text{old}}}(O|q)}}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\Big[r_{i,t}(\theta)\,R(q,o_{i}),\;\text{clip}\!\big(r_{i,t}(\theta),1-\epsilon,1+\epsilon\big)\,R(q,o_{i})\Big]\end{aligned}$    (2)
+AIME24 AIME25 AMC GSM8K MATH500 Minerva Olympiad Avg
+31.4 11.8 74.5 88.9 79.0 33.5 40.1 51.31 (-2.56)
+Entropy collapse
+$\begin{aligned} \mathcal{J}(\theta)&=\;\mathcal{J}_{GRPO}(\theta)+\alpha\;\mathbb{E}_{q\sim P(Q),\;\{o_{i}\}_{i=1}^{G}\sim\pi^{T}_{\theta_{\text{old}}}(O|q)}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\Big[r_{i,t}(\theta)\,{\color[rgb]{0.828125,0.18359375,0.18359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.828125,0.18359375,0.18359375}\hat{A}_{i,t}},\;\text{clip}\!\big(r_{i,t}(\theta),1-\epsilon,1+\epsilon\big)\,{\color[rgb]{0.828125,0.18359375,0.18359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.828125,0.18359375,0.18359375}\hat{A}_{i,t}}\Big]\end{aligned}$    (3)
+AIME24 AIME25 AMC GSM8K MATH500 Minerva Olympiad Avg
+32.7 11.0 73.2 88.1 79.0 35.7 41.2 51.56 (-2.31)
+Entropy collapse
 5 Experiments
 To validate the effectiveness of our methods, we present the experimental setup and results in the following sections.
-Our work is based on the EasyR1 and VeRL frameworks (Yaowei et al., 2025; Sheng et al., 2025), and we compare with the RL baselines GRPO (Shao et al., 2024) and its entropy-regularization variants (Hou et al., 2025; Cheng et al., 2025).
+Our work is based on the EasyR1 and VeRL frameworks (zheng2025easyr1; sheng2025hybridflow), and we compare with the RL baselines GRPO (shao2024deepseekmath) and its entropy-regularization variants (hou2025advancing; cheng2025reasoning).
 5.1 Experimental setup
-Model and Dataset: We conduct experiments to evaluate the effectiveness of AEPO in RFT for mathematical reasoning tasks.
-The base model is Qwen2.5-Math-7B (Yang et al., 2024), an LLM specialized for mathematical problem solving.
-For training, we use the DAPO-17K dataset (Yu et al., 2025), which contains diverse problem instances curated for RFT.
-Benchmark: Evaluation is performed on a broad suite of mathematical reasoning benchmarks, including AIME24 (HuggingFaceH4, 2025), AMC (Lightman et al., 2023), College Math (Tang et al., 2024), GSM8K (Cobbe et al., 2021), MATH (Lightman et al., 2023), Minerva Math (Lewkowycz et al., 2022), and Olympiad (Lightman et al., 2023).
+Model and Dataset: We conduct experiments to evaluate the effectiveness of AEPO in RL for mathematical reasoning tasks.
+The base models include Qwen2.5-7B, Qwen2.5-Math-7B and Qwen3-4B.
+(yang2024qwen2; yang2025qwen3).
+For training, we use the DAPO-17K dataset (yu2025dapo), which contains diverse problem instances curated for RL.
+Benchmark: Evaluation is performed on a broad suite of mathematical reasoning benchmarks, including AIME24, AIME25 (hf_aime2024), AMC (lightman2023lets), GSM8K (cobbe2021gsm8k), MATH (lightman2023lets), Minerva Math (lewkowycz2022solving), and Olympiad (lightman2023lets).
 These benchmarks collectively span a wide range of difficulty levels, from grade school arithmetic to advanced competition-level mathematics, and together they cover nearly all mainstream benchmarks for mathematical reasoning, enabling a comprehensive assessment of AEPO’s impact on reasoning performance across diverse tasks.
-Implementation: We initialize all policy models and conduct experiments on 8 A800 GPUs (40GB).
-Unless otherwise specified, we follow the default EasyR1 settings: maximum response length of 2048, global batch size 128, rollout batch size 512, rollout group size GG=5, temperature 1.0, learning rate $10^{-6}$, $\epsilon=0.2$, and a binary reward.
-For AEPO, $T_{\text{high}}=1.2$, $T_{\text{low}}=0.8$, we replace 60 temperature-adjusted positive samples in each batch.
-For entropy-regularized baselines, we set $\lambda=0.03$ for Entropy-Reg, and $\beta=0.4$, $\kappa=2$ for Entropy-Adv (Cheng et al., 2025).
-To exclude potential confounding factors, we did not apply KL divergence in any of our experiments.
 5.2 Main results
-Table 2 summarizes the performance of AEPO compared with GRPO and entropy-based baselines across seven mathematical reasoning benchmarks.
-We observe that AEPO consistently outperforms all competing methods, achieving the best results on every benchmark.
-In particular, AEPO yields an average score of 61.36, representing a relative improvement of +3.40 points over GRPO.
-Notably, the gains are especially pronounced on high-difficulty datasets such as AIME24 and AMC, where exploration plays a critical role in finding correct reasoning paths.
-Compared to entropy-regularized variants, AEPO also exhibits clear advantages.
-Both Entropy-Reg and Entropy-Adv deliver only marginal improvements over GRPO.
-As shown in Fig.
-4, the entropy regularization term in Entropy-Reg introduces significant bias: during mid-stage training, entropy begins to surge uncontrollably, indicating that entropy gradually replaces accuracy as the dominant optimization signal, with accumulated bias distorting the learning process.
-Entropy-Adv, on the other hand, fails to fundamentally reverse the trend of entropy collapse, leaving the model with insufficient exploration.
-In contrast, AEPO achieves consistent improvements across all benchmarks, underscoring the benefits of principled entropy control over heuristic entropy bonuses.
-Figure 4: Entropy trajectories of AEPO compared with GRPO, Entropy-Reg, and Entropy-Adv.
-While GRPO and Entropy-Adv suffer from monotonic entropy collapse and fail to maintain exploration, Entropy-Reg leads to unstable fluctuations.
-In contrast, AEPO stabilizes entropy around the target level, demonstrating controllable and robust entropy regulation throughout training.
-5.3 Entropy and exploration
-Table 3: Comparison of different ablation losses: entropy control ability, and benchmark performance.
-Loss and performance Entropy control
-$\begin{aligned} \mathcal{J}(\theta)&=\;\mathcal{J}_{GRPO}(\theta)+\alpha\;\mathbb{E}_{q\sim P(Q),\;\{o_{i}\}_{i=1}^{G}\sim{\color[rgb]{0.828125,0.18359375,0.18359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.828125,0.18359375,0.18359375}\pi_{\theta_{\text{old}}}(O|q)}}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\Big[r_{i,t}(\theta)\,R(q,o_{i}),\;\text{clip}\!\big(r_{i,t}(\theta),1-\epsilon,1+\epsilon\big)\,R(q,o_{i})\Big]\end{aligned}$
-AIME24 AMC College GSM8K MATH Minerva Math Olympiad Avg
-33.3 75.0 47.7 88.9 80.0 33.5 40.1 56.93 (-4.43)
-entropy collapse
-$\begin{aligned} \mathcal{J}(\theta)&=\;\mathcal{J}_{GRPO}(\theta)+\alpha\;\mathbb{E}_{q\sim P(Q),\;\{o_{i}\}_{i=1}^{G}\sim\pi^{T}_{\theta_{\text{old}}}(O|q)}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\Big[r_{i,t}(\theta)\,{\color[rgb]{0.828125,0.18359375,0.18359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.828125,0.18359375,0.18359375}\hat{A}_{i,t}},\;\text{clip}\!\big(r_{i,t}(\theta),1-\epsilon,1+\epsilon\big)\,{\color[rgb]{0.828125,0.18359375,0.18359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.828125,0.18359375,0.18359375}\hat{A}_{i,t}}\Big]\end{aligned}$
-AIME24 AMC College GSM8K MATH Minerva Math Olympiad Avg
-36.7 75.0 48.3 88.1 80.8 36.0 41.8 58.14 (-3.22)
-entropy collapse
-To further examine the relationship between entropy and exploration, we evaluate AEPO under different entropy thresholds $\mathcal{H}$.
-As shown in Table 2, AEPO consistently surpasses GRPO across all benchmarks regardless of the entropy target, with strong improvements on challenging datasets such as AIME24 and MATH.
-At different thresholds, AEPO maintains clear advantages, indicating that principled entropy regulation yields benefits across a wide range of entropy levels.
-Figure 1 further highlights this effect.
-While GRPO exhibits monotonic entropy collapse, AEPO stably maintains entropy around arbitrary preset levels simply by adjusting the hyperparameter $\mathcal{H}$, since AEPO is not sensitive to the parameter $\alpha$, as long as it is chosen within a reasonable range.
-This controllability allows us to probe the direct impact of entropy on exploration and downstream performance.
-Notably, we observe a non-monotonic trend: performance first improves as entropy rises from low to moderate levels, but begins to decline when entropy becomes excessively high.
-This provides explicit evidence that moderate entropy fosters effective exploration, while overly high entropy disperses optimization and harms accuracy.
-Together, these results confirm that AEPO not only eliminates entropy collapse but also clarifies the link between entropy, exploration, and performance, offering a controllable pathway to balance exploration with convergence.
+Breaking the entropy bottleneck.
+Fig.
+1 and 4 show that AEPO fundamentally resolves the entropy collapse issue that hampers GRPO and other RL baselines: AEPO maintains entropy around arbitrary target levels and avoids both collapse and uncontrolled expansion.
+As summarized in Table 1, on almost all mathematical reasoning benchmarks, AEPO surpasses GRPO and entropy-based baselines, demonstrating a much stronger capacity for effective exploration.
+In contrast, entropy-regularized baselines either fail to sustain entropy or introduce significant optimization bias that disrupts learning.
+Breaking the exploration bottleneck.
+AEPO delivers consistent improvements over RL baselines on both pass@1 and pass@$k$ across every benchmark.
+More importantly, as shown in the Table.
+2, AEPO shows the ability to explore broader than the base model on pass@1024, demonstrating that RL can meaningfully expand the reasoning frontier rather than merely sharpening the base model’s existing knowledge.
+During training, the pass@1024 performance of GRPO decreases monotonically, whereas AEPO shows the ability to expand the reasoning frontier, exhibiting upward trends rather than a purely decreasing curve.
+These results confirm that exploration can genuinely unlock new reasoning behaviors that are inaccessible to standard RL methods when guided by principled entropy modulation.
+Entropy-performance relation.
+By varying the entropy target, AEPO enables a controlled study of how entropy affects exploration and downstream accuracy.
+Across all entropy levels, AEPO consistently outperforms GRPO, demonstrating that stable entropy regulation is inherently beneficial.
+More interestingly, as shown in Table 3, AEPO reveals a clear non-monotonic relationship: moderate entropy improves exploration and enhances performance, whereas overly high entropy disperses optimization and reduces accuracy.
+Different benchmarks also exhibit varying degrees of benefit from exploration; some continue to gain as entropy increases, while others experience performance degradation.
+Overall, the results display a rise-then-fall trend, indicating the presence of an optimal entropy regime for effective reasoning.
 6 Ablation Study
-To assess the contribution of each component in AEPO, we design two ablation studies to verify the necessity of distribution as regularization and REINFORCE as regularization for achieving effective entropy control.
-6.1 Distribution as regularization
-One critical component of AEPO is the use of temperature-adjusted distributions for entropy control.
+To assess the contribution of each component in AEPO, we design two ablation studies to verify the necessity of temperature as regularization and REINFORCE as regularization for achieving effective entropy control.
+6.1 Temperature as regularization
+One critical component of AEPO is the use of temperature-adjusted samples for entropy control.
 In our design, the REINFORCE regularization term samples from a modified distribution, where the temperature $T$ is adaptively adjusted based on the previous step’s entropy.
 This adjustment ensures that positive samples carry either higher or lower entropy as required, thereby stabilizing the overall entropy around the target threshold $\mathcal{H}$.
-To validate the necessity of this design, we replace the temperature-adjusted distribution with the original distribution, as shown in Table 3.
+To validate the necessity of this design, we replace the temperature-adjusted distribution with the original distribution, as shown in Eq.
+(3) of Table 3.
 The results demonstrate that when REINFORCE samples are drawn directly from the original policy distribution (i.e., without temperature adjustment), entropy control collapses: the policy entropy monotonically decreases during training, similar to GRPO.
-More importantly, the average benchmark score drops to 57.39, which is even worse than standard GRPO and far below AEPO (61.36).
-This shows that the variance of REINFORCE gradients, when unregularized by distribution adjustment, further degrades optimization stability and downstream reasoning accuracy.
-These findings confirm that distribution adjustment is indispensable for AEPO: it directly enables controllable entropy regulation, reduces variance, and consistently improves reasoning performance across benchmarks.
-Moreover, it provides strong evidence for the role of policy gradient as regularization: the REINFORCE term in AEPO influences the exploration ability of the GRPO component through entropy control, thereby shaping the overall optimization dynamics during training.
+More importantly, the average benchmark score drops to 51.31, which is even worse than standard GRPO and far below AEPO (53.87).
+This shows that the variance of REINFORCE gradients, when unregularized by distribution adjustment, further degrades optimization performance.
+These findings confirm that temperature adjustment is indispensable for AEPO: it directly enables controllable entropy regulation, and consistently improves reasoning performance across benchmarks.
+Moreover, it provides strong evidence for the relation between entropy and exploration: the REINFORCE term in AEPO influences the exploration ability of the GRPO component through entropy control, thereby shaping the overall optimization dynamics during training.
 6.2 REINFORCE as regularization
@@
 To examine the role of REINFORCE, we conduct an ablation where the regularization term is still sampled from temperature-adjusted distributions, but the advantage function $\hat{A}_{t}$ is used directly without filtering negative samples.
-As shown in Table 3, this variant fails to prevent entropy collapse, leading to degraded performance across benchmarks.
-The average score drops to 58.14.
+As shown in Eq.
+(3) of Table 3, this variant fails to prevent entropy collapse, leading to degraded performance across benchmarks.
+The average score drops to 51.56.
 This confirms that filtering negative samples via REINFORCE is indispensable.
 7 Conclusion and discussion
-In this paper, we propose Arbitrary Entropy Policy Optimization (AEPO), a principled RFT framework that addresses one of the most persistent challenges in RFT—precise and stable entropy control.
+In this paper, we propose AEPO, a principled RL framework that addresses one of the most persistent challenges in RL—precise and stable entropy control.
 Unlike traditional entropy regularization methods that trade off exploration against stability, AEPO achieves controllable entropy regulation through a unified design that integrates policy gradient, distribution, and REINFORCE as regularization components.
 This formulation eliminates the entropy collapse phenomenon in GRPO and maintains policy entropy within an arbitrarily specified range, enabling balanced and consistent exploration throughout training.
-Extensive experiments across seven mathematical reasoning benchmarks demonstrate that AEPO consistently outperforms entropy-based baselines, exhibiting greater stability, generalization, and robustness to hyperparameters.
+Extensive experiments across seven mathematical reasoning benchmarks demonstrate that AEPO consistently outperforms entropy-based baselines on both pass@1 and pass@k, exhibiting greater stability, generalization, and robustness to hyperparameters.
+Moreover, AEPO shows that RL-based exploration can indeed move beyond the limitations of the base model, confirming that reinforcement learning is capable of expanding the reasoning frontier rather than merely refining pretrained knowledge.
 More importantly, AEPO reveals a non-monotonic relationship between entropy and reasoning performance, showing that moderate entropy fosters exploration while excessive entropy impairs optimization—offering the first quantitative evidence linking entropy dynamics to reasoning capability in large language models.
-Beyond solving entropy collapse, AEPO establishes a generalizable paradigm for learning under target distributions, suggesting its applicability to broader domains such as multi-modal alignment and long-horizon reasoning.
-In summary, AEPO transforms entropy from a passive statistical indicator into an active, tunable variable that fundamentally governs exploration, stability, and performance in RFT.
+Beyond entropy control, AEPO provides a broadly generalizable framework for learning under target distributions.
+The temperature-based mechanism used in AEPO should be viewed only as a special case: temperature creates the target distribution.
+When the task objective changes, the same principle naturally extends beyond entropy control.
+For example, if the goal is to mitigate overthinking by encouraging shorter responses, and if one can construct a transformed distribution $\pi^{*}$ from the current policy $\pi$ that produces more concise outputs while preserving correctness, then AEPO can use samples from $\pi^{*}$ to form a REINFORCE regularization.
+In doing so, the current policy is gradually guided toward acquiring the desired properties of $\pi^{*}$.
+More generally, whenever a target distribution with specific behavioral characteristics can be derived from the current policy, AEPO draws samples from this target and applies a REINFORCE-based regularization term, enabling the current policy to absorb the desired behavioral characteristics from the target distribution.
+AEPO is not merely an entropy-control technique but a general paradigm for steering policies toward arbitrarily defined target behaviors through principled policy-gradient regularization.
+The same mechanism that adjusts the entropy can therefore be extended to new domains, illustrating the strong generalizability of AEPO.
 Declaration of AI
 AI is only used for translation and language polishing in this paper.
+Appendix A Detailed implementation:
+We follow the default EasyR1 setup for all experiments and run all models on 8 A800 GPUs. The full training configuration is listed in Table 4.
+Table 4: Detailed implementation for all experiments.
+[…]
+RL Settings:
+   Maximum response length 8192
+   Batch size 512
+   Mini batch size 128
+   Rollout group size $G$ 8
+   Sampling temperature 1.0
+   Learning rate $1\times 10^{-6}$
+   Clip range $\epsilon=\epsilon_{low}=\epsilon_{high}$ 0.2
+   Reward type Binary reward
+   AEPO Settings:
+    High temperature $T_{\text{high}}$ 1.2
+    Low temperature $T_{\text{low}}$ 0.8
+    Temperature-adjusted REINFORCE samples 60 (positive samples per step) for entropy up
+30 for entropy down
+   Entropy-Regularized Baselines:
+    Entropy-Reg coefficient $\lambda$ 0.015
+    Entropy-Adv parameters $(\beta,\kappa)$ (0.4, 2)
+Evaluation Settings:
+   Maximum response length 8192
+   Top P 0.95
+   Temperature 0.1 for Pass@1
+1.0 for Pass@k
+Appendix B Proof of Theorem 4.3
+Reinforcement Learning. Let $o$ denote a response sampled from the policy $\pi_{\theta}(\cdot|s_{0})$ for query $s_{0}$. Given a reward function $R(s_{0},o)\in\mathbb{R}$, the policy objective is
+$J(\theta)=\mathbb{E}_{s_{0}\sim\rho}\mathbb{E}_{o\sim\pi_{\theta}(O|s_{0})}[R(o)]$
+$\nabla_{\theta}J(\theta)=\mathbb{E}_{s_{0}\sim\rho}\mathbb{E}_{o\sim\pi_{\theta}(O|s_{0})}[R(o)\cdot\sum_{t=0}^{T-1}\nabla_{\theta}log\pi_{\theta}(o_{t}|s_{t})]$
+Definition B.1.
+policy entropy: $H(\theta)=-\mathbb{E}_{s_{0}\sim\rho}\mathbb{E}_{o\sim\pi_{\theta}(O|s_{0})}[\frac{1}{|o|}\sum_{t=0}^{|o|-1}\mathbb{E}_{o_{t}\sim\pi_{\theta}(\cdot|s_{t})}[\text{log}\pi_{\theta}(o_{t}|s_{t})]]$
+To simplify the analysis, let us consider the particular case in which $|o|=1$. For a given initial state $s$, a corresponding action space $\mathcal{A}$, and a given reward function
+$R(s,a)=\begin{cases}1\quad\text{if }\ a\in\mathcal{A}^{*}\\ 0\quad\text{otherwise}\end{cases}$
+where $\mathcal{A}^{*}$ denotes the set of reference actions for a given state $s$, we have some simplified definitions as follows:
+Definition B.2.
+simplified objective: $J(\theta)=\mathbb{E}_{a\sim\pi_{\theta}(\cdot|s)}[R(s,a)]$
+Definition B.3.
+simplified policy entropy: $H(\theta)=-\mathbb{E}_{a\sim\pi_{\theta}(\cdot|s)}[\text{log}\pi_{\theta}(a|s)]$
+We first consider an intrinsic property of LLMs: they are softmax policies, which means the policies are parameterized by
+$\pi_{\theta}(a|s)=\frac{\text{exp}(z(s,a))}{\sum_{a^{\prime}\in\mathcal{A}}\text{exp}(z(s,a^{\prime}))}$
+where $z(s,a)$ is the logit for the state-action pair $(s,a)$ under parameter $\theta$. Furthermore, given a temperature T, we define:
+$\pi_{\theta}^{T}(a|s)=\frac{\text{exp}(\frac{z(s,a)}{T})}{\sum_{a^{\prime}\in\mathcal{A}}\text{exp}(\frac{z(s,a^{\prime})}{T})}$
+we write $z_{\theta_{k}}(s,a)$ simply as $z_{k}(s,a)$, and we write $\pi_{\theta}^{T}(a|s)$ simply as $\pi_{\theta}(a|s)$ when $T=1$.
+Problem 1.
+To optimize the RL objective above, we usually apply the gradient ascent method. We can calculate the policy gradient as
+$\nabla_{\theta}J(\theta)=\mathbb{E}_{a\sim\pi_{\theta}(\cdot|s)}[R(s,a)\cdot\nabla_{\theta}\text{log}\pi_{\theta}(a|s)],$
+And then we can update the parameter according to the following rule:
+$\theta_{k+1}=\theta_{k}+\eta\cdot\nabla_{\theta}J(\theta_{k})=\theta_{k}+\eta\cdot\mathbb{E}_{{\color[rgb]{0.828125,0.18359375,0.18359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.828125,0.18359375,0.18359375}a\sim\pi_{\theta_{k}}(\cdot|s)}}[R(s,a)\cdot\nabla_{\theta}\text{log}\pi_{\theta_{k}}(a|s)].$
+Empirically, we observe that applying the aforementioned update rule leads to a decreasing trend in policy entropy.
+This leads to the natural question: How is this trend affected if we modify the sampling distribution for policy gradient estimation?
+We consider a more general parameter update rule based on the standard gradient ascent rule.
+$\theta_{k+1}=\theta_{k}+\eta\cdot\mathbb{E}_{{\color[rgb]{0.828125,0.18359375,0.18359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.828125,0.18359375,0.18359375}a\sim\pi_{\theta_{k}}^{T}(\cdot|s)}}[R(s,a)\cdot\nabla_{\theta}\text{log}\pi_{\theta_{k}}(a|s)]$ (4)
+When we set the temperature $T=1$, the update rule above becomes identical to the standard gradient ascent method.
+In the following, we will investigate how the policy entropy changes with the parameter update defined as Eq.4
+Assumption B.4.
+$\nabla_{\theta}z_{k}(s,a),\nabla_{\theta}z_{k}(s,b)>=c_{a,b}\cdot\delta_{a,b}$, $\forall a,b\in\mathcal{A}$.
+where $c_{a,b}\in\mathbb{R}$ and $\delta_{a,b}=\textbf{1}[a=b]$.
+In the following, we set $c_{a,b}\equiv 1$.
+Lemma B.5.
+Let the sequence of parameters $\{\theta_{k}\}$ be governed by the update rule described above. Denote the function
+$\Delta H_{k}(T)=-\eta\cdot\sum_{a^{*}\in\mathcal{A}^{*}}\pi_{\theta_{k}}^{T}(a^{*}|s)\cdot\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a^{*}}-\pi_{\theta_{k}}(a^{\prime}|s)).$
+Then we have
+$\Delta H_{k}(T)\approx H(\theta_{k+1})-H(\theta_{k}),$
+and
+$\Delta H_{k}^{{}^{\prime}}(T)|_{T=1}>0\quad\text{if}\quad\Delta H_{k}(T)|_{T=1}<0.$
+Proof.
+For a given parameter $\theta_{k}$ and a relatively small learning rate $\eta$, leveraging Taylor’expansion under first-order approximation, we have
+$H(\theta_{k+1})-H(\theta_{k})\approx<\nabla_{\theta}H(\theta_{k}),\theta_{k+1}-\theta_{k}>$
+We then to derive what $\nabla_{\theta}H(\theta_{k})$ is, according to the definition of $H$, we have
+$\nabla_{\theta}H(\theta_{k})$ $=-\nabla_{\theta}\mathbb{E}_{a\sim\pi_{\theta_{k}}(\cdot|s)}[\text{log}\pi_{\theta_{k}}(a|s)]$
+$=-\nabla_{\theta}\sum_{a\in\mathcal{A}}[\pi_{\theta_{k}}(a|s)\cdot\text{log}\pi_{\theta_{k}}(a|s)]$
+$=-\mathbb{E}_{a\sim\pi_{\theta_{k}}}[\nabla_{\theta}\text{log}\pi_{\theta_{k}}(a|s)\cdot\text{log}\pi_{\theta_{k}}(a|s)+\nabla_{\theta}\text{log}\pi_{\theta_{k}}(a|s)]$
+$=-\mathbb{E}_{a\sim\pi_{\theta_{k}}}[\nabla_{\theta}\text{log}\pi_{\theta_{k}}(a|s)\cdot\text{log}\pi_{\theta_{k}}(a|s)]$
+the last equality follows from $\sum_{a\in\mathcal{A}}\nabla_{\theta}\pi_{\theta_{k}}(a|s)=\nabla_{\theta}\sum_{a\in\mathcal{A}}\pi_{\theta_{k}}(a|s)=\nabla_{\theta}1=0$.
+Then we have
+$H(\theta_{k+1})-H(\theta_{k})$ $\approx-<\mathbb{E}_{a\sim\pi_{\theta_{k}}(\cdot|s)}[\nabla_{\theta}\text{log}\pi_{\theta_{k}}(a|s)\cdot\text{log}\pi_{\theta_{k}}(a|s)],\theta_{k+1}-\theta_{k}>$
+$=-\mathbb{E}_{a\sim\pi_{\theta_{k}}(\cdot|s)}[\text{log}\pi_{\theta_{k}}(a|s)\cdot<\nabla_{\theta}\text{log}\pi_{\theta_{k}}(a|s),\theta_{k+1}-\theta_{k}>]$
+$\approx-\mathbb{E}_{a\sim\pi_{\theta_{k}}(\cdot|s)}[\text{log}\pi_{\theta_{k}}(a|s)\cdot(\text{log}\pi_{\theta_{k+1}}(a|s)-\text{log}\pi_{\theta_{k}}(a|s))]$
+$\approx-\mathbb{E}_{a\sim\pi_{\theta_{k}}(\cdot|s)}[\text{log}\pi_{\theta_{k}}(a|s)\cdot\sum_{a^{\prime}\in\mathcal{A}}\frac{\textbf{d}\text{log}\pi_{\theta_{k}}(a|s)}{\textbf{d}z_{k}(s,a^{\prime})}\cdot\Delta z_{k}(s,a^{\prime})]$
+We now evaluate the two terms, $\frac{\textbf{d}\text{log}\pi_{\theta_{k}}(a|s)}{\textbf{d}z_{k}(s,a^{\prime})}$ and $\Delta z_{k}(s,a^{\prime})$, in turn.
+$\frac{\textbf{d}\text{log}\pi_{\theta_{k}}(a|s)}{\textbf{d}z_{k}(s,a^{\prime})}=\delta_{a,a^{\prime}}-\pi_{\theta_{k}}(a^{\prime}|s);$
+$\Delta z_{k}(s,a^{\prime})$ $\approx<\frac{\textbf{d}z_{k}(s,a^{\prime})}{\textbf{d}\theta},\theta_{k+1}-\theta_{k}>$ (5)
+$=<\nabla_{\theta}z_{k}(s,a^{\prime}),\eta\cdot\mathbb{E}_{{\color[rgb]{0.828125,0.18359375,0.18359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.828125,0.18359375,0.18359375}a\sim\pi_{\theta_{k}}^{T}(\cdot|s)}}[R(s,a)\cdot\nabla_{\theta}\text{log}\pi_{\theta_{k}}(a|s)]>$
+$=\eta\cdot\sum_{a^{*}\in\mathcal{A}^{*}}\pi_{\theta_{k}}^{T}(a^{*}|s)\cdot<\nabla_{\theta}z_{k}(s,a^{\prime}),\nabla_{\theta}z_{k}(s,a^{*})-\mathbb{E}_{a\sim\pi_{\theta_{k}}}[\nabla_{\theta}z_{k}(s,a)]>$
+$=\eta\cdot\sum_{a^{*}\in\mathcal{A}^{*}}[\pi_{\theta_{k}}^{T}(a^{*}|s)\cdot(\delta_{a^{\prime},a^{*}}-\pi_{\theta_{k}}(a^{\prime}|s))]$
+where the first equality in Eq.5 follows from
+$\nabla_{\theta}\text{log}\pi_{\theta_{k}}(a|s)$ $=\nabla_{\theta}[z_{k}(s,a)-\text{log}\sum_{a^{\prime}\in\mathcal{A}}\text{exp}(z_{k}(s,a^{\prime}))]$
+$=\nabla_{\theta}z_{k}(s,a)-\frac{1}{\sum_{a^{\prime}\in\mathcal{A}}\text{exp}(z_{k}(s,a^{\prime}))}(\sum_{a^{\prime}\in\mathcal{A}}\text{exp}(z_{k}(s,a^{\prime}))\cdot\nabla_{\theta}z_{k}(s,a^{\prime}))$
+$=\nabla_{\theta}z_{k}(s,a)-\mathbb{E}_{a\sim\pi_{\theta_{k}}}[\nabla_{\theta}z_{k}(s,a)]$
+and the last equality in Eq.5 follows from assumption 1. Then, we have
+$H(\theta_{k+1})-H(\theta_{k})$ $\approx-\sum_{a^{\prime}\in\mathcal{A}}\Delta z_{k}(s,a^{\prime})\cdot\mathbb{E}_{a\sim\pi_{\theta_{k}}(\cdot|s)}[\text{log}\pi_{\theta_{k}}(a|s)\cdot\frac{\textbf{d}\text{log}\pi_{\theta_{k}}(a|s)}{\textbf{d}z_{k}(s,a^{\prime})}]$ (6)
+$=-\sum_{a^{\prime}\in\mathcal{A}}\Delta z_{k}(s,a^{\prime})\cdot\mathbb{E}_{a\sim\pi_{\theta_{k}}(\cdot|s)}[\text{log}\pi_{\theta_{k}}(a|s)\cdot(\delta_{a,a^{\prime}}-\pi_{\theta_{k}}(a^{\prime}|s))]$
+$=-\sum_{a^{\prime}\in\mathcal{A}}\Delta z_{k}(s,a^{\prime})\cdot[\pi_{\theta_{k}}(a^{\prime}|s)\cdot\text{log}\pi_{\theta_{k}}(a^{\prime}|s)-\pi_{\theta_{k}}(a^{\prime}|s)\cdot\mathbb{E}_{a\sim\pi_{\theta_{k}}(\cdot|s)}[\text{log}\pi_{\theta_{k}}(a|s)]]$
+$=-(\mathbb{E}_{a\sim\pi_{\theta_{k}}(\cdot|s)}[\Delta z_{k}(s,a)\cdot\text{log}\pi_{\theta_{k}}(a|s)]-\mathbb{E}_{a\sim\pi_{\theta_{k}}(\cdot|s)}[\Delta z_{k}(s,a)]\cdot\mathbb{E}_{a\sim\pi_{\theta_{k}}(\cdot|s)}[\text{log}\pi_{\theta_{k}}(a|s)])$
+$=-\text{Cov}_{\pi_{\theta_{k}}}(\Delta z_{k}(s,a),\text{log}\pi_{\theta_{k}}(a|s))$
+$=-\text{Cov}_{\pi_{\theta_{k}}}(\Delta z_{k}(s,a),z_{k}(a|s)),$
+take Eq.5 into Eq.6, we have
+$H(\theta_{k+1})-H(\theta_{k})$ $\approx-\eta\cdot\sum_{a^{*}\in\mathcal{A}^{*}}\pi_{\theta_{k}}^{T}(a^{*}|s)\cdot\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a),\delta_{a,a^{*}}-\pi_{\theta_{k}}(a|s))\equiv\Delta H_{k}(T)$
+We now proceed to the proof of the second part of the lemma. We first partition the set $\mathcal{A}^{*}$ into three subsets: $\mathcal{A}_{1}$, $\mathcal{A}_{2}$, and $\mathcal{A}_{3}$, where
+$\mathcal{A}_{1}=\{a|a\in\mathcal{A}^{*},z_{k}(s,a)<\mathbb{E}_{a^{\prime}\sim\pi_{\theta_{k}}(\cdot|s)}[z(s,a^{\prime})]\};$
+$\mathcal{A}_{2}=\{a|a\in\mathcal{A}^{*},z_{k}(s,a)\geq\mathbb{E}_{a^{\prime}\sim\pi_{\theta_{k}}(\cdot|s)}[z(s,a^{\prime})],\text{Cov}_{a^{\prime}\sim\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a})<C_{\pi_{k}}\};$
+$\mathcal{A}_{3}=\{a|a\in\mathcal{A}^{*},z_{k}(s,a)\geq\mathbb{E}_{a^{\prime}\sim\pi_{\theta_{k}}(\cdot|s)}[z(s,a^{\prime})],\text{Cov}_{a^{\prime}\sim\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a})\geq C_{\pi_{k}}\}$
+where $C_{\pi_{k}}=\text{Cov}_{a^{\prime}\sim\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\pi_{\theta_{k}}(a^{\prime}|s))$. According to the condition, we have
+$0$ $\leq\sum_{a^{*}\in\mathcal{A}^{*}}\pi_{\theta_{k}}(a^{*}|s)\cdot\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a),\delta_{a,a^{*}}-\pi_{\theta_{k}}(a|s))$
+$=\sum_{i=1}^{3}\sum_{a^{*}\in\mathcal{A}_{i}}\pi_{\theta_{k}}(a^{*}|s)\cdot\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a),\delta_{a,a^{*}}-\pi_{\theta_{k}}(a|s))$
+$\leq\sum_{a^{*}\in\mathcal{A}_{2}\bigcup\mathcal{A}_{3}}\pi_{\theta_{k}}(a^{*}|s)\cdot\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a),\delta_{a,a^{*}}-\pi_{\theta_{k}}(a|s))$
+We now turn to the analysis of $\frac{\textbf{d}\Delta H_{k}(T)}{\textbf{d}T}$.
+$\frac{\textbf{d}\pi_{\theta_{k}}^{T}(a|s)}{\textbf{d}T}$ $=\frac{1}{(\sum_{a^{\prime}\in\mathcal{A}}\text{exp}(\frac{z_{k}(s,a^{\prime})}{T}))^{2}}\cdot[\text{exp}(\frac{z_{k}(s,a)}{T})(-T^{-2})z_{k}(s,a)\sum_{a^{\prime}\in\mathcal{A}}\text{exp}(\frac{z_{k}(s,a^{\prime})}{T})$
+$-\text{exp}(\frac{z_{k}(s,a)}{T})\sum_{a^{\prime}\in\mathcal{A}}(-T^{-2})\text{exp}(\frac{z_{k}(s,a^{\prime})}{T})z_{k}(s,a^{\prime})]$
+$=-T^{-2}\cdot\pi_{\theta_{k}}^{T}(a|s)\cdot[z_{k}(s,a)-\mathbb{E}_{\pi_{\theta_{k}}^{T}}[z_{k}(s,a^{\prime})]]$
+when we take $T=1$, we have
+$\frac{\textbf{d}\pi_{\theta_{k}}^{T}(a|s)}{\textbf{d}T}|_{T=1}=-\pi_{\theta_{k}}(a|s)\cdot[z_{k}(s,a)-\mathbb{E}_{\pi_{\theta_{k}}}[z_{k}(s,a^{\prime})]]=-\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a,a^{\prime}})$
+$\frac{\textbf{d}\Delta H(T)}{\textbf{d}T}|_{T=1}$ $=-\eta\cdot\sum_{a^{*}\in\mathcal{A}^{*}}\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a^{*}}-\pi_{\theta_{k}}(a^{\prime}|s))\cdot\frac{\textbf{d}\pi_{\theta_{k}}^{T}(a^{*}|s)}{\textbf{d}T}$
+$=-\eta\cdot\sum_{a^{*}\in\mathcal{A}^{*}}\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a^{*}}-\pi_{\theta_{k}}(a^{\prime}|s))\cdot(-\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a^{*}}))$
+$=\eta\cdot[\sum_{a^{*}\in\mathcal{A}_{1}\bigcup\mathcal{A}_{2}\bigcup\mathcal{A}_{3}}\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a^{*}}-\pi_{\theta_{k}}(a^{\prime}|s))\cdot\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a^{*}})]$
+We now proceed with a case-by-case analysis based on different $a^{*}$.
+For $a^{*}\in\mathcal{A}_{1}$, we have $z_{k}(s,a)<\mathbb{E}_{a^{\prime}\sim\pi_{\theta_{k}}(\cdot|s)}[z(s,a^{\prime})]$, and then we can get $\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a^{*}})<0$, also we have $\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a^{*}}-\pi_{\theta_{k}}(a^{\prime}|s))<0$, therefore
+$\sum_{a^{*}\in\mathcal{A}_{1}}\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a^{*}}-\pi_{\theta_{k}}(a^{\prime}|s))\cdot\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a^{*}})\geq 0,$
+For $a^{*}\in\mathcal{A}_{2}$, we have $z_{k}(s,a^{*})\geq\mathbb{E}_{a^{\prime}\sim\pi_{\theta_{k}}(\cdot|s)}[z(s,a^{\prime})]$ and $\text{Cov}_{a^{\prime}\sim\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a^{*}})<C_{\pi_{k}}$, we set $u=\max_{a^{*}\in\mathcal{A}_{2}}\{z_{k}(s,a^{*})\}$, therefore
+$\sum_{a^{*}\in\mathcal{A}_{2}}\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a^{*}}-\pi_{\theta_{k}}(a^{\prime}|s))\cdot\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a^{*}})$
+$=$ $\sum_{a^{*}\in\mathcal{A}_{2}}\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a^{*}}-\pi_{\theta_{k}}(a^{\prime}|s))\cdot\pi_{\theta_{k}}(a^{*}|s)\cdot[z_{k}(s,a^{*})-\mathbb{E}_{\pi_{\theta_{k}}}[z_{k}(s,a^{\prime})]]$
+$=$ $\sum_{a^{*}\in\mathcal{A}_{2}}\underbrace{\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a^{*}}-\pi_{\theta_{k}}(a^{\prime}|s))}_{<0}\cdot\pi_{\theta_{k}}(a^{*}|s)\cdot[\underbrace{z_{k}(s,a^{*})-u}_{\leq 0}+u-\mathbb{E}_{\pi_{\theta_{k}}}[z_{k}(s,a^{\prime})]]$
+$\geq$ $\sum_{a^{*}\in\mathcal{A}_{2}}\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a^{*}}-\pi_{\theta_{k}}(a^{\prime}|s))\cdot\pi_{\theta_{k}}(a^{*}|s)\cdot[u-\mathbb{E}_{\pi_{\theta_{k}}}[z_{k}(s,a^{\prime})]],$
+For $a^{*}\in\mathcal{A}_{3}$, we have
+$\sum_{a^{*}\in\mathcal{A}_{3}}\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a^{*}}-\pi_{\theta_{k}}(a^{\prime}|s))\cdot\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a^{*}})$
+$=$ $\sum_{a^{*}\in\mathcal{A}_{3}}\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a^{*}}-\pi_{\theta_{k}}(a^{\prime}|s))\cdot\pi_{\theta_{k}}(a^{*}|s)\cdot[z_{k}(s,a^{*})-\mathbb{E}_{\pi_{\theta_{k}}}[z_{k}(s,a^{\prime})]]$
+$=$ $\sum_{a^{*}\in\mathcal{A}_{3}}\underbrace{\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a^{*}}-\pi_{\theta_{k}}(a^{\prime}|s))}_{>0}\cdot\pi_{\theta_{k}}(a^{*}|s)\cdot[\underbrace{z_{k}(s,a^{*})-u}_{\geq 0}+u-\mathbb{E}_{\pi_{\theta_{k}}}[z_{k}(s,a^{\prime})]]$
+$\geq$ $\sum_{a^{*}\in\mathcal{A}_{3}}\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a^{*}}-\pi_{\theta_{k}}(a^{\prime}|s))\cdot\pi_{\theta_{k}}(a^{*}|s)\cdot[u-\mathbb{E}_{\pi_{\theta_{k}}}[z_{k}(s,a^{\prime})]],$
+Therefore, combining the results from the three cases discussed above, we arrive at the following conclusion:
+$\frac{\textbf{d}\Delta H(T)}{\textbf{d}T}|_{T=1}$
+$=$ $\eta\cdot[\sum_{a^{*}\in\mathcal{A}_{1}\bigcup\mathcal{A}_{2}\bigcup\mathcal{A}_{3}}\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a^{*}}-\pi_{\theta_{k}}(a^{\prime}|s))\cdot\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a^{*}})]$
+$\geq$ $\eta\cdot[\sum_{a^{*}\in\mathcal{A}_{2}\bigcup\mathcal{A}_{3}}\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a^{*}}-\pi_{\theta_{k}}(a^{\prime}|s))\cdot\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a^{*}})]$
+$\geq$ $\eta\cdot[\sum_{a^{*}\in\mathcal{A}_{2}\bigcup\mathcal{A}_{3}}\pi_{\theta_{k}}(a^{*}|s)\cdot\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a^{*}}-\pi_{\theta_{k}}(a^{\prime}|s))\cdot(u-\mathbb{E}_{\pi_{\theta_{k}}}[z_{k}(s,a^{\prime})])]$
+$=$ $\eta\cdot(u-\mathbb{E}_{\pi_{\theta_{k}}}[z_{k}(s,a^{\prime})])\cdot[\sum_{a^{*}\in\mathcal{A}_{2}\bigcup\mathcal{A}_{3}}\pi_{\theta_{k}}(a^{*}|s)\cdot\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a^{*}}-\pi_{\theta_{k}}(a^{\prime}|s))]$
+$\geq$ $0$